### v2

# Axion Icebergs: Clockwork ALPs at hadron colliders

Abstract
Scenarios with multiple pseudoscalars are interesting as they usually tend to provide a framework to naturally realize a light axion with a large decay constant which has rich applications in cosmology, especially in the context of inflation and light dark matter physics. On the other hand, from a particle physics perspective, this facilitates a solution to the strong CP problem with a low Peccei-Quinn symmetry breaking scale. One such realization is afforded within the framework of the clockwork mechanism where the axion can have suppressed couplings with the gluons or photons while its companion axion-like particles (ALPs) have relatively unsuppressed couplings, thereby facilitating detectability. We study a minimal clockwork model for the QCD axion invoking a KSVZ-like setup and examine the visibility of its unique multi-ALP $(a_{n})$ signature at the LHC, the most sensitive channel being $pp\to a_{n}\,(+\,{\rm additional\,jets})$ followed by $a_{n}\to\gamma\gamma$. In congruence with the astrophysical and cosmological bounds for the axion, a striking feature emerges for the case of light ALPs $(m\sim\mathcal{O}(10\,{\rm GeV}))$ wherein the mass-splittings among the former are so small that the signal profile mimics that of a single broad resonance, or an axion iceberg. The scenario is found to be imminently testable by the end of LHC’s Run 3 phase for an integrated luminosity of $\sim 300{\rm\,fb^{-1}}$. A larger average ALP mass, on the other hand, results in multiple closely-spaced peaks with a characteristic signal profile, and would be expected to be seen at the forthcoming HL-LHC. Possible additional signals are also listed.
preprint: APS/123-QED
I Introduction
The strong CP problem is, perhaps, one of the strongest motivations to look for new dynamics beyond the Standard Model (SM) of particle physics. Of particular interest has been establishing the existence of the light pseudoscalar particle — popularly called an axion or, more specifically, the QCD axion — predicted by the Peccei-Quinn (PQ) mechanism Peccei and Quinn (1977); Weinberg (1978); Wilczek (1978) which is, arguably, the simplest and the most elegant solution [^1] There exist interesting alternatives to the PQ scenario, most notable among them being solutions based on a massless up quark Georgi and McArthur (1981); Kaplan and Manohar (1986); Choi et al. (1988); Banks et al. (1994); Dine et al. (2015) and those invoking spontaneous CP violation (a la the Nelson-Barr mechanism) Nelson (1984); Barr (1984). However, the $m_{u}=0$ solution within the standard theory currently stands in tension with the lattice QCD result of the topological mass contribution Alexandrou et al. (2020). Similarly, models based on the Nelson-Barr mechanism are also subject to several theoretical challenges Dine and Draper (2015). to the problem proposed to date.
Some theoretical and cosmological issues such as the quality problem and domain wall formation do question the mechanism’s general viability, although there have been several noteworthy developments towards mitigating these issues Lazarides and Shafi (1982); Chatterjee et al. (2020); Ibe et al. (2020); Zhang (2024). Nevertheless, the axion solution offers an interesting and important domain to explore phenomenologically, more so because the axion can potentially also play the role of an ultralight dark matter candidate Duffy and van Bibber (2009); Chadha-Day et al. (2022).
The axion is typically expected to have a very large decay constant ($f\gtrsim\mathcal{O}(10^{10}\,\mbox{GeV})$) —with the lower limit depending on its mass—so as to be consistent with the current limits (drawn from astrophysical and cosmological observations as well as from direct search experiments) on its effective couplings with photons (as also electrons and nucleons) Navas et al. (2024); O’Hare (2020). In other words, this essentially mandates that the PQ symmetry breaking scale ($f_{PQ}$) must be hierarchically larger than the electroweak (EW) scale. Furthermore, the QCD axion mass typically scales inversely with the decay constant and, hence, a large $f$ also implies a very light mass for the axion. The extremely large value of $f_{PQ}$ (and, the attendant smallness of the axion mass), thus, introduces a new hierarchy problem.
In addition, the tiny mass of the minimal QCD axion and its extremely feeble couplings render it nearly invisible to most of the current and future experimental probes, especially in the context of high energy experiments. Consequently, recent collider searches have tended to concentrate, instead, either on axion-like particles (ALPs), that have no role in the strong CP issue Bauer et al. (2019); d’Enterria (2021); Flórez et al. (2021), or on heavy QCD axions obtained by way of introducing new or extended confining sectors Rubakov (1997); Berezhiani et al. (2001); Fukuda et al. (2015); Hook et al. (2020); Agrawal and Howe (2018); Gherghetta et al. (2020) The latter case, although interesting in its own right, typically does not capture the cosmological attributes of a light axion, i.e. a role as a possible dark matter candidate and/or as the inflaton field in the so-called natural inflation scenario Freese et al. (1990); Adams et al. (1993). In view of this, a question naturally arises as to whether the discovery of a heavy axion would necessarily negate the existence of the light QCD axion. This aspect assumes particular significance as there are hints of several local excesses in searches for such scalar/pseudoscalar resonances Aad et al. (2023a); CMS (2023); Aad et al. (2022); Sirunyan et al. (2019); Aad et al. (2024a); Schmieden (2021) without these being robust enough to merit a discovery claim.
Scenarios with multiple aligned axions Kim et al. (2005); Choi et al. (2014) could, in principle, address these issues without assuming a large hierarchy in the individual decay constants and yet allowing for a quasi-flat direction in the field space, thereby leading to an effective decay constant $f_{a}\gg f_{PQ}$ for the QCD axion. Ref.Higaki et al. (2016a) sought to exploit this with $f_{PQ}$ close to the EW scale and ALPs with masses $\gtrsim m_{EW}$.
In this work, we seek to present a theoretically better-motivated model that also leads to a very distinctive multi-ALP spectrum. With ALP masses $\lesssim m_{EW}$, this exposition would not only be complementary to the scenarios of refs.Higaki et al. (2016a); Farina et al. (2017) but would also confront features in existant data at the LHC Aad et al. (2023a); CMS (2023); Aad et al. (2022); Sirunyan et al. (2019); Aad et al. (2024a); Schmieden (2021). Furthermore, the light QCD axion inherent to the scenario, apart from being a viable DM candidate, would likely be testable at the current and/or next generation of dedicated experiments. To this end, we focus on the clockwork paradigm, originally proposed in refs.Choi et al. (2014); Choi and Im (2016); Kaplan and Rattazzi (2016); Giudice and McCullough (2017), as it automatically generates hierarchical couplings or mass scales through the localization of the lightest particle in the theory space lattice defined by $N$ fields which interact with each other through nearest-neighbour couplings with a strength characterized by an ${\cal O}(1)$ parameter $q$. The massive modes, on the other hand, are delocalized over the entire lattice.
To realize a QCD axion within such a construction, we consider a model with a clockwork sector comprising $N+1$ complex scalars $\Phi_{j}(j=0,...,N)$, each charged under a unique $U(1)$ symmetry. Inspired by the KSVZ model Kim (1979); Shifman et al. (1980), we introduce a pair of new, $SU(2)_{L}$ singlet, coloured Weyl fermions ($\Psi_{L,R}$), chiral under the $N$th group $U(1)_{N}$, which couple to the $N$-th complex scalar. With only a single species of new coloured fermions, the putative domain wall problem, on account of a spontaneous breaking of the discrete symmetry present in the QCD-induced potential for the axion, is well under control, and such a KSVZ-like model is free of the problems that often beset, say, a DFSZ-like Zhitnitsky (1980); Dine et al. (1981) scenario.
A spontaneous breaking of the full symmetry ($U(1)^{N+1}$) at a scale $f$ (an analogue of $f_{PQ}$) in the clockwork sector, alongwith an explicit breaking of the symmetry to a single factor $U(1)_{CW}$ due to the nearest-neighbour interactions characterized by an effective mass scale $m\ll f$, leads to one massless Nambu-Goldstone Boson (NGB) and $N$ pseudo-NGBs in the spectrum. The chiral anomaly, then, induces a coupling of the light NGB to the gluonic topological operator ($\tilde{G}G$) with an exponentially large effective decay constant $\sim q^{N}f$, thanks to the localization mechanism. Below the confinement scale, this leads to a small mass for the light NGB through nonperturbative effects. Therefore, for $q>1$ and a sizable number of CW fields, the light NGB can be identified with the QCD axion even for the fundamental SSB scale $f$ being as low as a few TeVs. The pNGBs, on the other hand, act as ALPs with masses of the order $\sim mq$ (thereby creating a mass gap) and couple to gluons with an effective decay constant comparable to the scale $f$. The nonzero hypercharge of $\Psi_{L,R}$ not only allows them to decay, but also engenders loop-induced couplings of the ALPs to a pair of photons and $Z$’s. At a hadron collider (current or future), ALPs are dominantly produced through gluon-gluon fusion, and their most prominent, yet unique, signature would be the occurrence of multiple closely spaced resonance peaks in the $\gamma\gamma$, $ZZ$ and $Z\gamma$ channels. The closeness in the ALP masses emerges naturally in the model as a consequence of the clockwork construction wherein the mass-splitting varies as $\sim 2m/N$, (with $N\gtrsim 10$ for adequate enhancement of the axion decay constant). Understandably, for lighter ALPs ($m\lesssim 100$ GeV), the mass-splittings would be small, and in certain cases so much so that the currently operating detectors at the LHC might not be able to completely resolve the individual peaks, leading to the appearance of the entire ALP spectrum as a single broad resonance. We argue that both signals (multiple closely spaced narrow peaks or a single anomalously broad peak) are unique to the clockwork paradigm and can be generated neither by a single, phenomenologically viable ALP candidate, nor by a generic model of multiple ALPs without invoking an ad hoc fine-tuning in the parameter space. Additionally, such signatures are complementary to the possibility of observing periodic (in invariant mass) signals in the case of relatively heavier ALPs, e.g. in the spirit of the analyses carried out in the context of the gravity model in refs.Giudice et al. (2018); Beauchesne and Kats (2020); Aad et al. (2023b). Driven by this notion, we assume benchmark points with ALP masses at or below the EW scale and investigate their detection prospects in the diphoton channel at the LHC. The most interesting outcome is found in the case of the mass parameter $m$ being small, where although the individual resonances are barely visible over the background, the significance of the conjoint signal profile, composed of multiple overlapping peaks, is projected to reach the discovery threshold even as early as by the end of the LHC’s Run 3 phase. An analysis of the diphoton final state, which is arguably the cleanest channel to probe the ALPs, is also motivated by the number of dedicated collider searches for light scalar resonances, near and below the EW scale, performed by the ATLAS Aad et al. (2023a, c) and CMS CMS (2023) collaborations in the recent years. Thus, our study also appeals to the case for further experimental probes of light scalar or pseudoscalar particles in the diphoton channel, scanning not only for one but multiple closely-spaced candidates or even in the form of a broad resonance.
Note that although the introduction of a clockwork sector does not alter or improve upon the basic features of the PQ mechanism, it does provide ways to mitigate certain issues present in the minimal QCD axion models, which in turn makes this avenue even more desirable to explore. For one, the possibility of a low scale $f_{PQ}$ helps improve the quality of the axion solution against putative quantum gravity effects [^2] Since the degree of the global symmetry breaking depends on the exact manifestation of quantum gravity effects near the Planck scale, a robust estimate of the quality problem in our model, given the complexity of the global symmetry involved, is difficult to obtain. Nevertheless, under the assumption that the leading QG effect arises as a dim-5 operator of the form $(g/M_{Pl})\left(|\Phi_{j}|^{2}\right)^{2}\Phi_{j^{\prime}}$ in the IR, the required fine-tuning for an acceptable PQ solution can be characterized in terms of an upper limit on the coupling $|g|$, which, in a model with $f_{PQ}\sim\mathcal{O}(1\,\mbox{TeV})$, can be estimated to be $|g|\lesssim 10^{-12}/N^{4}$. E.g., $N\sim 30$ implies $|g|\lesssim 10^{-18}$. Although this seems a very stringent upper limit in the absolute sense, it is nevertheless many orders of magnitude larger than what is required in the minimal models with $f_{PQ}\sim 10^{12}$ GeV, namely $|g|\lesssim 10^{-55}$ Kamionkowski and March-Russell (1992). Thus, even with the aforementioned ambiguities, it can be qualitatively argued that the clockwork scenario offers some respite from the severe quality issues of the typical axion models, if not a solution. The severity may be relaxed further with the imposition of additional discrete symmetries in the UV in order to restrict lower dimensional operators. However, we desist from doing so. Kamionkowski and March-Russell (1992); Higaki et al. (2016a). Secondly, the geometric progression of the light axion’s distribution over the lattice can be exploited to obtain an enhanced coupling with photons, independent of the coupling to the gluons, thereby opening a window in the parameter space that falls outside of the usual QCD band on the photon-coupling vs axion-mass plane. This aspect was studied in detail in ref.Farina et al. (2017).
The structure of the paper is as follows. In section II, we define a minimal clockwork model for a KSVZ-like QCD axion and demonstrate how it addresses the strong CP problem. We identify, in section III, four benchmark points of the model suitable for probing the ALPs through the gluon-fusion channel at the LHC. In section IV, we discuss the corresponding signal and background profiles for $\sqrt{s}=13$ TeV and an integrated luminosity $\mathcal{L}=138\,{\rm fb}^{-1}$. Section V, on the other hand, briefly describes the implications of the vector-like quark and the heavy radial scalars present in the model. Finally, we summarize the results and conclude in section VI.
II The model
We begin by constructing a clockwork realization of the minimal QCD axion, adopting a KSVZ-like Kim (1979); Shifman et al. (1980) scenario. In the ensuing, we first detail the structure and spectrum of the clockwork sector and then proceed towards describing how the KSVZ setup can be implemented in the context of the clockwork mechanism.
II.1 Clockwork scalars
We define the clockwork (CW) sector in terms of $N+1$ complex scalars $\Phi_{j}$, each charged under a global Abelian group $U(1)_{j}$ described by the Lagrangian
$\begin{array}[]{rcl}{\cal L}_{\rm CW}&=&\displaystyle{\cal L}_{1}+{\cal L}_{2}\,,\\[6.45831pt] {\cal L}_{1}&\equiv&\displaystyle\sum_{j=0}^{N}\left[(\partial_{\mu}\Phi_{j}^{\dagger})(\partial^{\mu}\Phi_{j})-\lambda\left(\Phi_{j}^{\dagger}\Phi_{j}-\frac{f^{2}}{2}\right)^{2}\right]\,,\\[10.76385pt] {\cal L}_{2}&\equiv&\lambda^{\prime}\Lambda^{3-q}\sum_{j=0}^{N-1}\Phi_{j}^{\dagger}\Phi_{j+1}^{q}+\mbox{h.c.}\ ,\end{array}$ (1)
where $\lambda,\lambda^{\prime}$ are dimensionless real constants and $\Lambda\,(\ll f)$ is a characteristic scale associated with ${\cal L}_{2}$ (the exponent $q$ being, as yet, unspecified). To understand better the pieces ${\cal L}_{1,2}$, let us examine them in isolation. In the absence of ${\cal L}_{2}$ (i.e., in the limit $\lambda^{\prime}\to 0$), the Lagrangian has a $U(1)^{N+1}$ global symmetry, the spontaneous breaking of which, at the scale $f$ (through $\langle\Phi_{j}^{\dagger}\Phi_{j}\rangle=f^{2}/2$ for all $j$) would lead to $(N+1)$ Goldstone bosons $\pi_{j}$, with the complex scalars being representable as
$\Phi_{j}=\frac{1}{\sqrt{2}}(\phi_{j}+f)e^{i\pi_{j}/f}\ ,$ (2)
where $\phi_{j}$ are the corresponding massive scalars. Note that this breaking mechanism still retains the additional global discrete symmetry $\Phi_{j}\leftrightarrow\Phi_{k}$ inherent in ${\cal L}_{1}$.
The large global symmetry, discrete or continuous (and the attendant plethora of Goldstones in the broken phase) is neither well-motivated nor desirable. Clearly, it can be broken by arbitrary terms expressing interactions between the scalars. On the other hand, at least a single $U(1)$ needs to be present [^3] As previously mentioned, an exact global symmetry is unrealistic because even in the absence of new symmetry breaking effects one would expect the global $U(1)$ to be explicitly broken by quantum gravity effects. The Planck suppressed symmetry violating terms thus generated can, in general, upset the PQ solution through corrections to the QCD axion potential to be discussed later Kamionkowski and March-Russell (1992). While the small $f_{PQ}$ in our scenario already alleviates this so-called quality problem, it may be further assuaged by ensuring that the Planck suppressed operators in the theory appear only at a sufficiently high mass-dimension, e.g., by considering the CW scalars to be composites of entities belonging to new confining sector(s) Randall (1992); Coy et al. (2017). so that its spontaneous breaking could lead to the axion. The interaction Lagrangian ${\cal L}_{2}$ serves exactly this purpose [^4] The requirement $\Lambda\ll f$ at best introduces a little hierarchy and for $q=2$, as we would adopt, it represents only a soft breaking of the symmetry.. Representing perhaps the simplest set of operators which preserve a single $U(1)$ while exhibiting the clockwork mechanism, it can be understood in terms of nearest-neighbour interaction terms [^5] In general, one could also have non-nearest neighbour interactions while keeping the CW symmetry intact, an example being a term of the form $\Phi^{\dagger}_{j}\Phi^{q/p}_{j+1}...\Phi^{q^{p}/p}_{j+p}$. With reference to eq.(5) such a term adds, to the average mass scale of the pNGBs, a contribution $m_{a}^{(p)}\sim m(q^{p}/p)$ Ben-Dayan (2019). For the case $q=2$ (which would be our primary choice as a benchmark for the collider analysis to be discussed) renormalizability necessitates $p=1,2$ for which the overall pNGB mass scale changes only slightly with $m_{a}\sim\sqrt{2}m_{a}^{(1)}$. The case for $q=3$ is a little different with $m_{a}\sim 3m_{a}^{(1)}$, although the qualitative aspects of our results and conclusions in this work would be applicable just as well. that exhibit locality in a theory space and, hence, the lattice defined by the fields $\Phi_{j}$ (each specifying a site in the theory space) can be regarded as the fifth dimension [^6] A continuum theory that reproduces this theory on discretization, though, requires the introduction of certain additional features and we desist from discussing it. As for the exponent $q$ in ${\cal L}_{2}$, a non-integer value would imply compositeness, thereby adding a further layer of complications. The clockwork mechanism, which will be employed in the following discussions, necessitates $q>1$. On the other hand, for $q>3$ one would obtain nonrenormalizable operators in $\mathcal{L}_{CW}$ implying that the explicit symmetry breaking takes place due to new dynamics operational at the heavy scale $\Lambda\gg f$ and beyond. Furthermore, $q>3$ would also tend to destabilize the vacuum. To avoid such issues and for simplicity we consider a renormalizable theory with $1<q\leq 3$.
With the introduction of ${\cal L}_{2}$, the erstwhile discrete symmetry is completely broken, whereas the $U(1)^{N+1}$ symmetry breaks explicitly to one combination $U(1)_{CW}$ corresponding to the generator,
$\mathcal{Q}_{CW}=\sum_{j}\frac{\mathcal{Q}_{j}}{q^{j}},$ (3)
with $\mathcal{Q}_{j}$ being the generators (charges) corresponding to the individual $U(1)_{j}$’s. Consequently, only one combination of the $\pi_{j}$’s would now be a true Goldstone boson, while all the rest would gain masses much smaller than $f$. In other words, the latter are now only pNGBs. To be specific, the full spin-0 Lagrangian, in the broken phase, can be written as,
$\begin{array}[]{rcl}\mathcal{L}^{Full}_{\Phi}=\sum_{j=0}^{N}\Big{[}\frac{1}{2}\partial_{\mu}\phi_{j}\partial^{\mu}\phi_{j}-\frac{1}{4}\lambda\left(\phi_{j}+f\right)^{4}\Big{]}+2^{(1-q)/2}\lambda^{\prime}\Lambda^{3-q}\displaystyle\sum_{j=0}^{N-1}(\phi_{j}+f)(\phi_{j+1}+f)^{q}\cos{\frac{\pi_{j}-q\pi_{j+1}}{f}}\,.\end{array}$ (4)
The mass eigenvalues of the pseudoscalar system are given by
$m_{n}^{2}=\begin{cases}0\,&n=0\\ m^{2}\,\left[1+q^{2}-2q\cos{\frac{n\pi}{N+1}}\right]\,&n\neq 0\end{cases},$ (5)
where $m^{2}\equiv 2^{(1-q)/2}\lambda^{\prime}\Lambda^{3-q}f^{q-1}$ (thus, $m_{n}^{2}\ll f^{2}$ as was expected). The transformation relation between the unphysical basis $(\pi_{j})$ and the physical basis $(a_{n})$ is specified by a matrix $C$ (viz. $a_{n}\equiv\sum_{j=0}^{N}C_{nj}\pi_{j}$) with elements
$\begin{split}C_{0j}&=\mathcal{N}_{0}q^{-j}\,,\\ C_{nj}&=\mathcal{N}_{n}\left[q\sin{\frac{jn\pi}{N+1}}-\sin\frac{(j+1)n\pi}{N+1}\right].\end{split}$ (6)
Here $\mathcal{N}_{0}$ and $\mathcal{N}_{n}$ are the normalization factors
$\mathcal{N}_{0}=\sqrt{\frac{q^{2}-1}{q^{2}-q^{-2N}}}\ ,\qquad\quad\mathcal{N}_{n>0}=\frac{m}{m_{n}}\sqrt{\frac{2}{N+1}}\,.$ (7)
It is evident from eq.(6) that while the heavy pseudoscalars are nearly delocalized over the entire lattice, the massless mode — which will subsequently assume the role of the QCD axion $(a_{0})$ — is localized towards the $j=0$ site, i.e. it has an exponentially suppressed overlap with $\pi_{N}$ by virtue of the clockwork mechanism. Interestingly, this feature will eventually ensure that the light axion has a hierarchically large decay constant [^7] Note that the usage of the terminology “decay constant” here is different from that encountered in the context of processes such as $\pi^{\pm}\to l^{\pm}\nu$, where $\pi^{\pm}$ are the charged QCD pions. as compared to the heavy axions in the spectrum.
II.2 A KSVZ axion
Now that we have a massless pseudoscalar in the clockwork spectrum, we would like to promote it to the status of a realistic QCD axion. This is most easily carried out by embedding the clockwork sector in any one of the usual QCD axion models. Owing to the singlet nature of the CW scalars, the KSVZ construction is the simplest one. To this end, we introduce one generation of Weyl fermions $\Psi_{L},\Psi_{R}$ with the following nontrivial charges under the SM gauge group and the $N$-th (global) Abelian group [^8] The chiral configuration only at the $N$-site ensures that the color anomaly is localized at a single site which enables us to exploit the CW mechanism to generate a large effective decay constant for the axion, as we will see shortly. On the other hand, while it is possible to have the $\Psi_{L,R}$ charged nonchirally under all the additional $U(1)$s, this only adds a layer of complexity without any qualitative changes in the phenomenology. of the CW sector ($G_{SM}\times U(1)_{N}$),
Field $SU(3)_{c}$ $SU(2)_{L}$ $U(1)_{Y}$ $U(1)_{N}$
$\Phi_{N}$ 11 11 0 $\xi$
$\Psi_{L}$ 33 11 $Y_{\Psi}$ $\xi_{L}$
$\Psi_{R}$ 33 11 $Y_{\Psi}$ $\xi_{R}$
with the restriction that $\xi=\xi_{L}-\xi_{R}$. This is reminiscent of the chiral PQ charges typically assigned to the new (i.e., non-SM) fermions in KSVZ models. The SM fields are assumed to be uncharged under the new $U(1)$’s. For $\xi_{L}\neq\xi_{R}$, the new singlet quark cannot have a bare mass term. On the other hand, for the aforementioned choice of $\xi=\xi_{L}-\xi_{R}$, the new fermion may now interact with the CW sector through a Yukawa term localized at the $N$-th site of the lattice. However, with just this interaction, the new quark would be absolutely stable and, hence, phenomenologically untenable. The situation can be remedied only if it couples with the SM fermions which, for our charge assignments, can be most simply ensured when $\xi_{R}=0$ (and, thus, $\xi_{L}=\xi$) alongwith an appropriate choice of $Y_{\Psi}$. To be specific, we choose $Y_{\Psi}=2/3$, thereby allowing $\Psi$ to mix with the up-type quarks. The complete set of Yukawa interactions is, thus, given by
[…]
where $Q^{(\alpha)}_{L}$ are the SM quark doublets, $u_{R}^{(\alpha)}$ the up-type singlets and $H$ the SM Higgs doublet. As for the dimensionless couplings $\lambda_{\Psi}$, $y^{(\alpha)}_{\Psi}$ and $y^{{}^{\prime}(\alpha)}_{\Psi}$, the first determines the mass of the $\Psi$ while the other two determine its decay rate. For the sake of simplicity, we assume that $\Psi_{L,R}$ couple to only the third-generation quarks, thereby precluding sizable flavour-changing neutral currents involving the first two generations.
In addition, the CW sector can also couple with the SM Higgs through [^9] For simplicity, we assume uniform value of the coupling $\lambda_{\Phi-H}$ for all $j$. Relaxing this only adds layers of complications without any qualitative change.
$\mathcal{L}_{\Phi-H}=-\lambda_{\Phi H}\sum_{j=0}^{N}\Phi_{j}^{\dagger}\Phi_{j}H^{\dagger}H\,.$ (9)
We will come back to reviewing the dynamics of the heavy radial scalars and the quarks, alongwith the flavour constraints on the Yukawa couplings, in a later section. Alluding to the core objective of the study, we first describe the low-energy physics of the pseudoscalars in the discussions below.
After SSB in the clockwork sector, we have
$\mathcal{L}_{\Psi}\supset-\frac{1}{\sqrt{2}}\lambda_{\Psi}(\phi_{N}+f)e^{i\xi\pi_{N}/f}\bar{\Psi}_{L}\Psi_{R}-y_{\Psi}\bar{Q}^{(3)}_{L}\tilde{H}\Psi_{R}-\frac{1}{\sqrt{2}}y^{\prime}_{\Psi}(\phi_{N}+f)e^{i\xi\pi_{N}/f}\bar{\Psi}_{L}u^{(3)}_{R}+\mbox{h.c.}\,.$ (10)
A rephasing of the new fermions with respect to the Goldstone field $\pi_{N}$ leads to the following interaction terms with the SM gauge bosons via the chiral anomaly Alonso-Álvarez et al. (2019) (see appendix A for a discussion)
$\mathcal{L}_{\pi VV}=-g_{\pi GG}\,\pi_{N}G^{A\mu\nu}\tilde{G}^{A}_{\mu\nu}-g_{\pi BB}\,\pi_{N}B^{\mu\nu}\tilde{B}_{\mu\nu},$ (11)
with the coefficients given by
$|g_{\pi GG}|=\frac{g^{2}_{s}}{32\pi^{2}f_{\rm eff}}\,,\quad|g_{\pi BB}|=\frac{2N_{c}g^{\prime 2}Y^{2}_{\Psi}}{32\pi^{2}f_{\rm eff}}\,$, (12)
and an effective scale defined through $f_{\rm eff}\,\equiv f/|\xi|$. Note that while $f$ continues to determine the masses of the $\Psi$ and the $\Phi_{N}$, it is $f_{\rm eff}$ that encapsulates the pNGB decay constants.
II.2.1 The Peccei-Quinn Mechanism
In order for the light pseudoscalar to be the quintessential QCD axion, it must offer a solution to the strong CP problem. Although ref.Farina et al. (2017) already has a brief discussion in this context, we outline here, for completeness, how the clockwork Lagrangian can be consistent with the usual Peccei-Quinn mechanism. The foremost requirement for the mechanism is the presence of a global symmetry (at least an accidental one) in the theory. In our case, this symmetry would correspond to the residual clockwork symmetry described in the preceding section (see eq.(3)). Now, in eq.(11), we can perform a shift in the $\pi_{N}$ field so as to exactly cancel the CP violating topological term $\bar{\theta}(\alpha_{S}/8\pi)\tilde{G}^{\mu\nu}G_{\mu\nu}$. This, however, is a true cancellation, and a potential solution to the strong CP problem, only when $\pi_{N}$ does not acquire a nonzero vacuum expectation value (VEV) after employing the constant shift. It is important to note that the explicit breaking of the $U(1)]^{N+1}$ symmetry in the clockwork sector, courtesy the nearest-neighbour interactions, does not generate a nonzero VEV for the pseudoscalar fields thanks to the residual shift symmetry $U(1)_{CW}$ Banerjee et al. (2018). The only other source from where $\pi_{N}$ may acquire a nontrivial potential are nonperturbative effects below the QCD confinement scale, leading to an effective potential of the form Di Vecchia and Veneziano (1980); Grilli di Cortona et al. (2016)
$V(\pi_{N})=-m_{\pi}^{2}f_{\pi}^{2}\sqrt{1-\frac{4m_{u}m_{d}}{(m_{u}+m_{d})^{2}}\sin^{2}{\left(\frac{\pi_{N}}{2f_{\rm eff}}\right)}}$ (13)
where $f_{\pi}$ and $m_{\pi}$ are the pion decay constant and mass, respectively, and $m_{u,d}$ are the masses of the light SM quarks $u$ and $d$. This shows that $\pi_{N}$ does not, indeed, acquire a nonzero VEV [^10] Naively, the potential would have a degenerate set of minima at $\langle\pi_{N}\rangle=2n\pi f_{\rm eff}$ for integral values of $n$. However, such nonzero VEVs are unphysical as they can always be shifted to zero by the residual symmetry transformation $\pi_{N}\to\pi_{N}+2n\pi f_{\rm eff}$ as a result of the $2\pi$ periodicity of the potential.. Expanding the potential in powers of $\pi_{N}$, the quadratic term engenders a mass correction to $\pi_{N}$
$\Delta m_{\pi_{N}}^{2}=\frac{f_{\pi}^{2}m_{\pi}^{2}}{f_{\rm eff}^{2}}\frac{m_{u}m_{d}}{(m_{u}+m_{d})^{2}}\,.$ (14)
Going to the physical basis of the CW axions, this generates a small mass for the lightest state, i.e., the KSVZ QCD axion, given by
$m_{a_{0}}\simeq\frac{f_{\pi}m_{\pi}}{f_{\rm eff}\,q^{N}}\frac{\sqrt{m_{u}m_{d}}}{(m_{u}+m_{d})},$ (15)
[…]
Figure 1: Limits on the effective decay constant $f_{0}$ of the light QCD axion vs its mass. Plot adapted from ref.O’Hare (2020).
It is perhaps prudent to remark at this juncture about the possible nature and implications of the topological defects that may arise in the multi-axion configuration of the clockwork sector. The spontaneous breaking of the $U(1)^{N+1}$ symmetry at a temperature $T\sim f\gg\Lambda_{QCD}$ (post inflation and reheating) is expected to lead to the formation of $N+1$ cosmic strings bound to one another through domain walls Long (2018); Higaki et al. (2016b); Lee et al. (2024). Such a string-wall network would, in principle, tend to radiate light axions as well as ALPs (as also gravitational waves Higaki et al. (2016b); Lee et al. (2024)) during the course of its cosmological evolution towards the epoch of the QCD phase transition [^11] The ALPs, owing to their relatively small decay constant, decay promptly once they are produced.. In a canonical framework, these emissions are estimated to scale as $f^{2}$ and, therefore, constitute a large fraction of the axion abundance in the usual QCD axion cosmology where $f\sim f_{0}\gtrsim 10^{9}\,\mbox{GeV}$. In the present context, though, the clockwork mechanism generates an exponential separation between the SSB scale $f$ and the effective decay constant of the light axion $f_{0}$, such contributions to the axion abundance is exceedingly suppressed for $f\sim\mathcal{O}(\mbox{TeV})$. At the epoch of the QCD phase transition, the induced potential of eq.(13) defines a unique vacuum for the system associated with a domain wall number $N_{DW}=1$ — a direct consequence of invoking only a single generation of the extra $SU(2)$ singlet coloured fermion $\Psi$. Similar to the case of a single axion Barr et al. (1987); Kim (1987); Chang et al. (1999), such a potential destabilizes the string-wall network and leads to its eventual collapse [^12] Thus, apart from the obvious simplicity, the choice of a single species of the singlet heavy quark possesses the important feature of alleviating the domain wall problem Sikivie (1982) typical to axion models., with its energy being released primarily in the form of axions and ALPs. This contribution to the axion abundance, too, can be estimated to be insignificant, with the axion emission being suppressed by the factor $\sim(m_{a_{0}}/m)^{3}$ Long (2018).
While a certain abundance of the axion as dark matter is expected to be generated via radiation and collapse of the topological defects in the clockwork system, a robust estimate of the same can only be obtained from a dedicated numerical simulation which is extremely complicated to perform for a multi-axion setup, especially for $N\gg\mathcal{O}(1)$. In the absence of such analyses, we admit to a reasonable assertion, based on the aforementioned arguments, that for the SSB scale of interest $f\sim\mathcal{O}(\mbox{TeV})\ll f_{0}$ the dominant contribution to the axion DM abundance emanates from the misalignment mechanism Fox et al. (2004), namely
$\Omega h^{2}\simeq 0.2\left(\frac{f_{0}}{10^{11}{\rm GeV}}\right)^{7/6}\left(\frac{\theta_{0}^{2}}{\pi^{2}/3}\right)$ (16)
where $\theta_{0}$ is the average initial misalignment angle of the QCD axion. This implies that for the natural choice of $\theta_{0}\sim\mathcal{O}(1)$ in eq.(16), the constraint from overclosure demands $f_{0}\lesssim 10^{11}\,\mbox{GeV}$. The experimental and observational limits demand, on the other hand, $f_{0}\gtrsim 10^{9}\,\mbox{GeV}$. The allowed window is then wide enough and comparable with the one usually encountered for the traditional QCD axion. Indeed, this is why it is also the target region for several cavity haloscope experiments such as ADMX, CAPP, HAYSTAC, QUAX, etc. (see e.g. figure 90.5 of the PDG review on axions Navas et al. (2024)).
In the spirit of these arguments, we would be assuming, in the discussions to follow, benchmarks that are consistent with the above mentioned bounds on $f_{0}$.
II.3 Axion physics above the QCD confinement scale
Concerned with the physics observable at high energy colliders, it suffices to consider the theory at energy scales below the EW symmetry breaking scale but significantly above the QCD confinement scale. At these scales, the axion Lagrangian becomes
$\begin{split}\mathcal{L}_{\pi vv}=&-g_{\pi gg}\,\pi_{N}G^{A\mu\nu}\tilde{G}^{A}_{\mu\nu}-g_{\pi\gamma\gamma}\,\pi_{N}F^{\mu\nu}\tilde{F}_{\mu\nu}\\ &-g_{\pi\gamma Z}\,\pi_{N}F^{\mu\nu}\tilde{Z}_{\mu\nu}-g_{\pi ZZ}\,\pi_{N}Z^{\mu\nu}\tilde{Z}_{\mu\nu}.\end{split}$ (17)
The coefficients in this case are given by
[…]
where $s_{w}\equiv\sin\theta_{w}$ and $c_{w}\equiv\cos\theta_{w}$ with $\theta_{w}$ being the Weinberg angle. For the mass-eigenstates, the couplings are simply scaled by the mixing parameters, viz.
$g^{(n)}_{agg}=g_{\pi gg}(f_{\rm eff}\to f_{n})$ (18)
where $f_{n}\equiv f_{\rm eff}/C_{nN}$ denotes the effective decay constant of the $n$-th physical axion. Using eq.(6), it is then readily apparent that the decay constants of the heavy pseudoscalars are hierarchically smaller than that of the light axion. For a SSB scale $f\sim\mathcal{O}(\mbox{TeV})$, this engenders significantly enhanced couplings of the heavy axions with gluons and the electroweak bosons as compared with those associated with typical light ALP candidates. This enhancement, in principle, can facilitate resonant production of the heavy CW axions at hadron colliders, which is going to be the main subject of discussion in the ensuing sections.
III Clockwork axions at colliders
As is already described above, the (light) axion, when compatible with the constraints from astrophysical and low-energy experiments Semertzidis and Youn (2022), is nearly invisible at high energy collider experiments. What seems promising, instead, is the prospect of probing the heavy pseudoscalars at the LHC and its forthcoming high-luminosity upgrade (HL-LHC). In this case, the production would be dominated by the gluon fusion channel with the hadronic cross-section given by
$\sigma(pp\to a_{n})=K^{0}_{\sigma}\frac{d\mathcal{L}_{gg}}{d\hat{s}}\Bigg{|}_{\hat{s}=m_{n}^{2}}\frac{\pi^{2}}{8m_{n}}\hat{\Gamma}^{(n)}_{gg}$, (19)
where $d\mathcal{L}_{gg}/d\hat{s}$ is the gluon-gluon luminosity and the $K$-factor $K^{0}_{\sigma}$ encapsulates the higher order QCD corrections. Using the MSTW2008nnlo68 parton densities, ref.Mariotti et al. (2018) estimates $K^{0}_{\sigma}\approx 3.7$. (including the full NNLO and approximate N3LO corrections Ball et al. (2013); Bonvini et al. (2014, 2016); Ahmed et al. (2016)) over the pseudoscalar mass range 4012540-125 GeV . In the absence of a full computation of the QCD corrections, we make the reasonable extrapolation that $K^{0}_{\sigma}\sim 3.7$ for the mass range 1012510-125 GeV. For ALP masses beyond $150$ GeV we assume a more conservative $K^{0}_{\sigma}=2.5$ Ahmed et al. (2016).
The leading (two-body) decay channels and the corresponding widths for the ALPs are
$\begin{split}\Gamma(a_{n}\to gg)&\equiv K_{gg}\hat{\Gamma}^{(n)}_{gg}=K_{gg}|C_{nN}g(x_{n})|^{2}\frac{\alpha_{S}^{2}}{32\pi^{3}}\frac{m_{n}^{3}}{f_{\mbox{eff}}^{2}}\\ \Gamma(a_{n}\to\gamma\gamma)&\equiv\Gamma^{(n)}_{\gamma\gamma}=9Y_{\Psi}^{4}|C_{nN}g(x_{n})|^{2}\frac{\alpha_{EM}^{2}}{64\pi^{3}}\frac{m_{n}^{3}}{f_{\mbox{eff}}^{2}}\\ \Gamma(a_{n}\to Z\gamma)&\equiv\Gamma^{(n)}_{Z\gamma}=9Y_{\Psi}^{4}|C_{nN}g(x_{n})|^{2}\frac{\alpha_{EM}^{2}\tan^{2}{\theta_{w}}}{8\pi^{3}}\frac{(m_{n}^{2}-m_{Z}^{2})^{3}}{m_{n}^{3}f_{\mbox{eff}}^{2}}\\ \Gamma(a_{n}\to ZZ)&\equiv\Gamma^{(n)}_{ZZ}=9Y_{\Psi}^{4}|C_{nN}g(x_{n})|^{2}\frac{\alpha_{EM}^{2}\tan^{4}{\theta_{w}}}{64\pi^{3}}\,\frac{(m_{n}^{2}-4m_{Z}^{2})^{3/2}}{f_{\mbox{eff}}^{2}}\end{split},$ (20)
where $x_{n}\equiv 4m_{\Psi}^{2}/m_{n}^{2}$ and the loop-integral $g(x)$ is defined as
$g(x)\equiv\left\{\begin{array}[]{lcl}\displaystyle x\left[\sin^{-1}(x^{-1/2})\right]^{2}&&x\geq 1\\[6.45831pt] \displaystyle\frac{-x}{4}\left[\ln\frac{1+\sqrt{1-x}}{1-\sqrt{1-x}}-i\pi\right]^{2}&&x<1\,.\end{array}\right$. (21)
In the expression for $\Gamma(a_{n}\to gg)$, the QCD correction amounts to $K_{gg}=2.1$ Djouadi (2008). Clearly, the ALPs decay dominantly to two gluons followed by decays to photons (and to $Z\gamma$ and $ZZ$ depending on kinematic feasibility). It is easy to infer—within the narrow width approximation— that the dijet cross-section would exceed the diphoton one by a factor of  [^13] Note that $K_{\gamma\gamma}\approx 1$. $2/(9Y_{\Psi}^{4})]K_{gg}(\alpha_{S}^{2}/\alpha_{EM}^{2})$. However, as far as detection at hadron colliders is concerned, the diphoton final state, understandably, offers better sensitivity as compared to the dijet channel. We will, therefore, consider only the processes $pp\to a_{n}\to\gamma\gamma$ for the case in hand.
The diphoton final state has been extensively studied in the literature in the context of spin-0 and spin-2 resonances, with dedicated searches by the ATLAS and CMS collaborations spanning almost the entire range of masses accessible with current sensitivities. With measurements thus far reporting, at best, only marginal excess of events over the SM background, the most recent analyses from ATLAS Aad et al. (2023a, 2021, 2024b) and CMS CMS (2023, 2024) place the strongest limits to date on the $pp\to\gamma\gamma$ fiducial cross-section for spin-0 resonances over a wide mass bracket ranging from nearly $10\,\mbox{GeV}$ to a few TeVs. Being optimistic that some new physics indeed exists beyond the SM, this broadly implies two possibilities —first, that the energy scale of new physics lies beyond the reach of the LHC and, second, that new dynamics exist within the LHC’s energy reach, albeit with feeble interaction strengths with the SM so as to be inaccessible with the current luminosity reach. In the following, we choose the latter viewpoint and explore viable scenarios where the diphoton cross-sections for the CW pseudoscalar spectrum could be very close to the current exclusion limits at the LHC obtained from the Run-2 data, such that a sizable enhancement in the signal significance can be achieved by the end of the ongoing Run-3 phase. This is interesting because even if a $5\sigma$ discovery looks improbable at the LHC, such enhancements could hint at a potential discovery at the HL-LHC taking into account the projected integrated luminosity of $\sim 3000\,\mbox{fb}^{-1}$, i.e. nearly 20 times the luminosity achieved by the end of LHC Run-2 ($\sim 138\,\mbox{fb}^{-1}$). We consider for our study three benchmark points categorized by the ALP masses [^14] While we do not consider ALP masses below $10$ GeV, collider searches for lighter axions in general have been studied in refs.Cid Vidal et al. (2019); Acanfora et al. (2024). in the range 103010-30 GeV, 3510535-105 GeV (with two sub-categories in this case) and 150450150-450 GeV, respectively, and have as guiding references the ATLAS analyses carried out in refs. Aad et al. (2023a, 2021, 2024b).
For each of the benchmark points, we choose $\xi=3$ and $q=2$ (as the middle point of the theoretically allowed range) and also assume $Y_{\Psi}=2/3$ so that the $\Psi$ may mix with the top quark. The other particulars are as follows—
Figure 2: (top) Masses and couplings for individual ALPs and corresponding diphoton cross-sections for benchmark I. The vertical axis on the right shows the extent of the field $\pi_{N}$ contained in the mass eigenstates $a_{n}$. (bottom) Mass-splittings between consecutive ALPs for benchmark I.
Benchmark I: We assume the parameter values $N=28$, $m=10$ GeV and $f=1000$ GeV in the CW sector. The resulting light axion in this case has a decay constant $f_{0}\approx 9\times 10^{10}$ GeV and a mass $m_{a_{0}}\approx 6.4\times 10^{-5}$ eV. The corresponding ALP masses span the range $\sim 10-30$ GeV. This particular choice of the CW parameters, especially the sizable number of particles $N$, is also motivated by the fact that it is compatible with the light axion being a DM candidate [^15] The notion of the KSVZ axion as dark matter has also been explored in ref.Alonso-Álvarez et al. (2023), albeit in the traditional one scalar setup. (with an acceptable relic density and consistency with cosmological observations)—as discussed in section II.2.1—for $\mathcal{O}(1)$ values of the misalignment angle [^16] See figures 4 and 5 in ref.Long (2018) for a summary of the various cosmological constraints on the CW parameter space.. With these choices in place, we can determine the diphoton cross-section in the narrow-width approximation (NWA). Fig.2 (top) shows the cross-sections for each of the ALPs (integrated over the full final state phase-space) computed for a center of mass energy $\sqrt{s}=13$ TeV using the NNPDF2.3LO PDF set Ball et al. (2015). Note that the cross-section peaks for an intermediate mode of the spectrum, the exact identity of which depends on the periodic variation of the couplings due to the clockwork mixing as well as the energy dependence of the parton-parton luminosity. Fig.2 (bottom), on the other hand, shows $\Delta m_{n}$ (the mass-splitting between consecutive CW modes) along the full spectrum, with the average mass-splitting being given by $\Delta m\sim 2m/N$. It is interesting to note that, with the consecutive differences $\Delta m_{n}\lesssim 1$ GeV over the entire extent of the ALP spectrum, the splittings are comparable with the prevailing detector resolution. Consequently, the individual resonances may not be entirely resolved and the events corresponding to the full spectrum (subject to the assumed bin size) may even appear as a single broad resonance [^17] This was also hinted at in ref.Higaki et al. (2016a). within an envelope of mass width  $\sim\,2m$ in the diphoton invariant mass distribution.
Benchmark II(a): The parameter configuration in the CW sector for this case is — $m=35$ GeV, $N=28$ and $f=1000$ GeV. The light axion has the same decay constant and mass as in benchmark I and, thus, is still a potential DM candidate, while the ALP masses now range from $35$ to $105$ GeV. The corresponding diphoton cross-sections are shown in Fig.3. Due to an increase in the value of the parameter $m$ (by a factor of 3.5) the mass-splittings in this case, readily inferred from eq.(5), are nearly triple in value compared to that in the preceding case. Therefore, in stark contrast with benchmark I, one expects to discern at least a few individual peaks in the invariant mass distribution of the signal events. We will see that this is indeed the case.
Benchmark II(b): Now, we assume a larger set of CW scalars, namely $N=40$, while keeping the rest of the parameters the same as in II(a). Then, from eq.(15) it is clear that the light axion mass lowers down to a value $m_{a_{0}}\sim 1.5\times 10^{-8}$ eV and the effective decay constant increases to $f_{0}\approx 4\times 10^{14}$ GeV. Note that an enhancement in $f_{0}$ of this magnitude is incompatible with the DM constraint unless $\theta_{0}\ll\mathcal{O}(1)$. Although certain cosmological scenarios could exist where a small initial misalignment angle seems plausible (see ref.Fox et al. (2004) for a discussion), we refrain from delving into those details and, from a strictly phenomenological standpoint of studying different ALP signatures at the colliders, assume a small enough value for $\theta_{0}$ such that the axion relic density is acceptable. The mass splittings in this case fall in the range $\sim 0.5-2.7$ GeV — smaller than what we obtained in benchmark II(a) but not small enough to be completely unresolved at the detector. Consequently, the distribution of signal events, as a function of the diphoton invariant mass is expected to be an undulating one.
Figure 3: Masses and couplings for individual ALPs and corresponding diphoton cross-sections. (top) benchmark II(a) and (bottom) benchmark II(b).
Benchmark III: Exploring ALPs heavier than the SM Higgs, we consider $m=150$ GeV, $N=40$ and $f=1600$ GeV. Due to the increase in the SSB scale $f$, the QCD axion mass is now even slightly lower than that in benchmark II(b), namely $m_{a_{0}}\approx 9.7\times 10^{-9}$ eV with $f_{0}\approx 6\times 10^{14}$ GeV. Naturally, the arguments pertaining to DM abundance as presented for benchmark II(b) are also applicable here. As before, Fig.4 shows the diphoton cross-sections for the ALP spectrum as well as the characteristic mass-splittings.
Figure 4: Diphoton cross-sections and mass-splittings for benchmark III.
With the model parameters defined, we now simulate the diphoton signal events and compare with the corresponding SM backgrounds for the $13$ TeV LHC. We write the Feynman rules of the models in Feynrules Alloul et al. (2014) and generate UFO files that are then imported in the MadGraph5 Alwall et al. (2014) event generator for a Monte Carlo simulation of the signal. Further, Pythia8 Sjöstrand et al. (2015); Bierlich et al. (2022) is used for parton showering and hadronization. Final state objects at the detector level are reconstructed using the fast simulation tool Delphes de Favereau et al. (2014). Within Delphes we choose to simulate the detector effects using the ATLAS detector card for the entirety our study.
The SM diphoton (with and without jets) background at the LHC has been very well-studied by both the ATLAS and CMS collaborations, not only in the context of the SM Higgs, but also for other exotic resonances, light and heavy. Indeed, there have been occasional reports of excesses, only to largely vanish on account of an even more careful recalibration of the backgrounds. We would be largely using the ATLAS analyses to estimate the backgrounds and, by extension, use the same kinematical restrictions etc. to estimate the signal strength as well.
IV Signal and Background profiles
While we would be concentrating on ALPs being produced in gluon fusion and decaying into a diphoton pair, note that the exact final state would be dependent on their average mass. In particular, a very light ALP, produced sans any accompanying high-$p_{T}$ particle, would result in a pair of relatively soft photons that would, typically, fail trigger requirements. In other words, a minimal number of high-$p_{T}$ entities must be present, and thus, both signal and background estimations would need to be done accordingly.
IV.1 Light ALPs
For Benchmark Point I, the ALP masses are in the $10-30\,\mbox{GeV}$ range, and the final state photons need to be boosted to overcome the detector’s trigger level energy threshold. To achieve this, we follow the strategy employed in the ATLAS search Aad et al. (2023a) (and also ref.Mariotti et al. (2018)) for low mass diphoton resonances and consider a diphoton final state with upto two additional jets. To simulate the signal events we first generate [^18] Understandably, this is overwhelmingly dominated by the subprocesses wherein the final state partons are gluons. $pp\to a_{n}+0/1/2\,{\rm partons}$ at the leading order (LO), using MadGraph5, for $\sqrt{s}=13$ TeV with the NNPDF2.3LO PDF set, followed by the two-body decay of the on-shell ALPs [^19] Given that the ALPs are spinless, there are no non-trivial spin correlations. Furthermore, given that they are very narrow, there is no real loss of information or accuracy in the neglect of possible off-shell effects, as can also be confirmed by the use of tools such as the Madspin Artoisenet et al. (2013) module.. This is followed by parton showering and hadronization via Pythia8 Sjöstrand et al. (2015); Bierlich et al. (2022). Matching and merging of the matrix elements (including the avoidance of overcounting) for showering is automated using the MLM matching scheme.
As mentioned previously, for a detector level simulation of the showered events we use the Delphes tool. The following summarizes the criteria used to identify and isolate final state objects.
Jets: Within Delphes, jets are reconstructed using the FastJet package Cacciari et al. (2012). Jet clustering is performed using the anti-$k_{T}$ algorithm Cacciari et al. (2008) with the jet cone radius parameter $R$ chosen to be 0.4. To be consonant with the ATLAS analyses, we require the minimum transverse momentum of jets ($p_{Tj}$) to be 20 GeV and its pseudorapidity to satisfy $|\eta_{j}|<2.5$. Finally, any two jets must be separated by $\Delta R(j,j)>0.7$ where $(\Delta R)\equiv\left[(\Delta\eta)^{2}+(\Delta\phi)^{2}\right]^{1/2}$ is the separation in the pseudorapidity-azimuthal angle plane.
Photons: For identification and isolation of photons, a cone size of $\Delta R=0.2$ around the photon candidate is considered. Denoting the ratio of the sum of transverse momenta of isolated objects (tracks, calorimeter towers, etc) to the candidate’s transverse momentum as $p_{T}^{\text{ratio}}$, it is demanded that $p_{T}^{\text{ratio}}<0.05$. Also incorporated is a $p_{T}$-dependent photon identification efficiency (following the ATLAS analysis of ref.Aad et al. (2023a)) that ranges from $\approx 70\%$ at $p_{T}=22$ GeV to $\approx 90\%$ for $p_{T}>50$ GeV.
In addition, we ensure angular separation between photons or jets by demanding that $\Delta R(\gamma,\gamma)>0.2$ and $\Delta R(\gamma,j)>0.4$. Finally, we demand events with at least two photons and at least a jet in the final state by imposing $N_{\gamma}\geq 2$ and $N_{j}\geq 1$. The above conditions are summarised in Table 1 for quick reference.
Channel : $pp\rightarrow a_{n}+0/1/2\,\text{jets}$   $a_{n}\rightarrow\gamma\gamma$
Acceptance Cuts
Photon identification
$\Delta R=0.2$, $p_{T}>0.5$ GeV, $p_{T}^{\text{ratio}}(\gamma)<0.05$
Jet ident ification
$\Delta R=0.4$ (anti-$k_{T}$), $p_{T}^{j}>20$ GeV, $|\eta_{j}|<2.5$
Isolation
$\Delta R(\gamma,\gamma)>0.2$, $\Delta R(\gamma,j)>0.4$, $\Delta R(j,j)>0.7$
$N_{\gamma}\geq 2$, $N_{j}\geq 1$
Table 1: Acceptance cuts for the final state objects in benchmarks I and II Aad et al. (2023a).
As for the QCD corrections to the process, within the narrow width approximation, we may factorise these separately for the production and the subsequent decay. For such masses of the ALPs, the QCD corrections to the production cross section can largely be summarised in terms of $K$-factors, with $K^{1}_{\sigma}\sim K^{2}_{\sigma}\simeq 2$ Gershtein et al. (2021) where the superscript indicates the number of jets in the final state. As for the decay, while the correction to the photonic branching fraction is small, that to the gluonic one is substantial, viz. $K_{gg}=2.1$ Djouadi (2008), and serves to scale the diphoton branching fraction (since $\mbox{BR}_{\gamma\gamma}\approx\Gamma_{\gamma\gamma}/\Gamma_{gg}$). For the effective $K$-factor, then, $K^{1,2}_{\sigma}/K_{gg}\approx 1$ for the two benchmark points with light ALPs.
Using MadAnalysis5 Conte et al. (2013) to analyze the signal events following the acceptance cuts, we order the photons and the jets (wherever applicable) in terms of their $p_{T}$ and present, in Fig.5 (top), the corresponding distributions, for an integrated luminosity of $\mathcal{L}=138\,\mbox{fb}^{-1}$. In this, we use a bin size of $1$ GeV, as in the ATLAS analysis Aad et al. (2023a). While the sharp edges (at 20 GeV) in the jet [^20] Note that the events with 2 jets are only a subset of the accepted set of events. $p_{T}$ distributions are but reflections of the acceptance cut imposed, the fast fall off at higher $p_{T}$s is char acteristic of QCD radiation. The positions of the peaks are just caused by an interplay of the two effects. As for the photons, the decay of an ALP, in its rest-frame, is isotropic with each daughter photon having an energy equalling half the ALP mass. This distribution, of course, has to be convoluted with the $p_{T}$ of the ALP itself (as counterbalancing the jet $p_{T}$s). And, finally, there are additional convolutions with both the ALP mass-spectrum and the corresponding production cross sections.
Figure 5: BP-I: event distributions, after applying acceptance cuts alone, vide. Table 1. (top) $p_{T}$ distributions for the leading photon ($\gamma_{1}$), sub-leading photon ($\gamma_{2}$), leading jet ($j_{1}$) and the sub-leading jet ($j_{2}$) and (bottom) the two-photon invariant mass.
Of particular interest is the diphoton invariant mass distribution (Fig.5 (bottom)) which encompasses the contributions of the entire spectrum of the $28$ ALPs. It is instructive to note the difference with the total production cross section (as in Fig.2 (top)) which shows a maximum for $m_{a}\approx 15$ GeV, owing, as we have discussed, to a combination of the variation in the coupling strength as well as the gluon-gluon flux. However, once the extra jets are required to be radiated, the kinematics (and the mass-dependence of the flux) does change considerably. Even more importantly, the requirements [^21] Further corrections arise from the $p_{T}$-dependence of the detector efficiencies. on $p_{T}(\gamma)$ serve to suppress the contribution from the very low mass ALPs, resulting in the significantly shifted maximum (now around $m_{\gamma\gamma}\approx 20$ GeV) as in Fig.5 (bottom).
For the background, diphoton (associated with additional jets)— initiated by $q\bar{q}$ as well as gggg— within the SM are considered. However, since there is a finite probability for a putative jet masquerading as a photon in the detector (this is especially true of a low energy jet depositing a large fraction of its energy in the electromagnetic calorimeter), one needs to consider such contributions as well. Thus, the backgrounds emanate primarily from
  • $pp\to\gamma\gamma$,
  • $pp\to\gamma j$,
  • $pp\to jj$.
in each case, accompanied by upto two additional jets.
Other backgrounds emanate from $e^{\pm}$ not leaving discernible tracks, thereby faking photons. These, though, contribute negligibly and can, therefore, be largely ignored. For an estimate of the relevant background distribution in the diphoton spectra, we again refer to the ATLAS analysis Aad et al. (2023a) wherein the fiducial region in the phase-space is defined using the selection cuts on $E^{\gamma}_{T}$, $p^{\gamma\gamma}_{T}$ and $\eta_{\gamma}$ as mentioned in Table 2. The diphoton background [^22] The background distribution as estimated in the ATLAS analysis Aad et al. (2023a) for a bin size of 1 GeV is available at https://atlas.web.cern.ch/Atlas/GROUPS/PHYSICS/PAPERS/HIGG-2019-23/. thus determined in the analysis has an uncertainty which is predominantly statistical.
Channel Event Selection Criteria
$N_{\gamma}=2$, $1\leq N_{j}\leq 2$,
$pp\rightarrow a_{n}+0/1/2\,\text{jets}$ $|\eta_{\gamma}|<2.37$ (excluding barrel-to-endcap region $1.37<|\eta_{\gamma}|<1.52$),
$a_{n}\rightarrow\gamma\gamma$ $E_{T}(\gamma)>$22 GeV, $p_{T}^{\gamma\gamma}>50$ GeV
Table 2: Selection cuts applied to form the fiducial signal regions for benchmarks I and II Aad et al. (2023a).
To examine the viability of the diphoton signal in the light of the ATLAS analysis, we must subject the former to the same selection criteria as the latter. These are listed in Table 2. As an examination of Fig.5 (top) shows, the $E_{T}(\gamma)$ cut, especially when applied to the subleading photon, does eliminate a non-negligible fraction of the signal events. Even harder is the cut on $p_{T}(\gamma\gamma)$, for this requires that the ALP (and, hence, the radiated off jets) must carry a substantial $p_{T}$. The consequent reduction of events is reflected in the $m_{\gamma\gamma}$ distribution as shown in Fig.6.
Figure 6: BP-I: Diphoton invariant mass distribution after applying selection cuts.
A striking aspect to note here (this was already evident in Fig.5 (bottom)) is that although the diphoton distribution receives contributions from all the ALPs in the spectrum, the separation between the individual resonances is smeared by the detector resolution [^23] We have made use of the energy resolution function for the ATLAS detector’s ECAL M. Aharrouche, et al, ATLAS Electromagnetic Barrel Calorimeter Collaboration (2006) as implemented within Delphes. For the phase-space region of interest throughout the presented analysis (i.e. for BP-I, II and III), the energy resolution of the individual photons varies roughly in the range $\sim[0.32,2.3]$ GeV. to a degree that the entire distribution appears as a single broad resonance — an iceberg of axions, so to say — with a FWHM $\sim 16$ GeV. This feature, as previously mentioned, is peculiar to the clockwork spectrum for any mass scale $m$ as long as $N$ is adequately large.
We can now estimate the significance of the signal events over the background bin-by-bin using $S=N_{sig}/\sqrt{N_{sig}+N_{bkg}}$, where $N_{sig}$ and $N_{bkg}$ denote the number of signal and background events, respectively, in a particular bin. Since the background uncertainty is statistics dominated, this gives quite a robust estimate of the sensitivity and can be bettered only by a dedicated search. The bin-wise significances (and the corresponding $p$-values) for $\mathcal{L}=138\,\mbox{fb}^{-1}$ with a bin width of $1$ GeV are depicted in Fig.7, and the maximum individual significance of $S^{1\,\rm GeV}_{138}\sim 1\sigma$ is obtained in the bin corresponding to $m_{\gamma\gamma}\sim 21-22$ GeV. It is tempting to increase the bin widths as this would be expected to substantially increase the per-bin significance owing to the individual peaks being closely packed within the invariant mass range 103010-30 GeV. However, since the existence of a wide (and finite) band of resonances is characteristic to the clockwork scenario, the shape of the invariant mass distribution could itself behave as a discriminator— more so in the case of small mass-splittings— and by enlarging the bin width one would, obviously, lose information about the underlying profile of the spectrum. In particular, this would help discriminate between a broad resonance and multiple sharp ones [^24] A further discriminant here would be the comparison between the width of the resonance and the total signal size. For a resonance, production cross section is simply related to the total width, whereas the size of the signal is just the product of the cross section, the branching fraction and the overall detector efficiency (the last being a known quantity).. A more useful method of estimating the total signal significance would be to consider the sum $S_{c}\equiv\sum_{i}S_{i}^{2}$ which, of course, is $\chi^{2}$-distributed for the appropriate degrees of freedom. Using this, we find for this benchmark point, the expected cumulative significance to be $S_{c,138}=3.38\sigma$ for $\mathcal{L}=138\,\rm fb^{-1}$. On the same note, it is worth alluding to the prospect that the significance can further build up by a sizable amount with an increase in the luminosity, even by as early as the end of the ongoing Run-3 phase of the LHC with a projected luminosity reach of $\sim 300\,\rm fb^{-1}$. A simple estimate of this can be obtained by scaling the current significance with the luminosity, i.e.
$S_{c,300}=S_{c,138}\times\sqrt{\frac{300}{138}}\approx 4.98\sigma\,$, (22)
which makes this benchmark point imminently testable by the end of Run 3.
Apropos of the current benchmark point, Fig.8 shows the variation of $S_{c}$ with the SSB scale $f$ for the $138$ $\rm fb^{-1}$ luminosity case alongwith a projection for the $300$ $\rm fb^{-1}$ reach of the LHC. The variation follows the approximate relation $S_{c}\propto f^{-2}$ which can be understood from the fact that, for a particular choice of the CW parameters, the pertinent cross-sections scale roughly as $f^{-2}$ with the pole masses being quite insensitive to it.
Figure 7: Bin-wise significance for benchmark I at the LHC for $\mathcal{L}=138\,\rm fb^{-1}$.
Figure 8: Cumulative significance as a function of the SSB scale $f$ for benchmark-I at the Run 2 (red curve) and Run 3 LHC (blue curve).
At this stage, it is worth examining if such a signal could have emanated from any other theory. First and foremost, the decay of the resonance into a diphoton pair rules out a spin-one particle. While spin-two or higher are possible, such fundamental particles are disfavoured on theoretical grounds (e.g., a graviton, while well-motivated, would not only be much narrower, but would also decay democratically into other channels). Such arguments narrow down the possibilities to only a scalar or a pseudoscalar. This apart, a width comparable to the mass (as is the case here), would require the couplings to be in the nonperturbative regime Salvio et al. (2016); Franceschini et al. (2016). (For instance, a single ALP resonance with a total decay width and mass $\Gamma(a)\sim m_{a}\sim\mathcal{O}(10{\rm\,GeV})$ would require its decay constant [^25] The inequality takes into account the dependence on the constituents of the anomaly loop. to be $f\lesssim 100$ MeV!). Irrespective of the underlying theory, such a large width would have translated, vide eq.(19), to a total production cross section several orders of magnitude larger than what is the case here. This would, immediately, raise question as to the reason for the suppression of the diphoton branching ratio that such a production cross section (alongwith the observed rates) would imply. Furthermore, an entity that couples so strongly would leave tell-tale signatures in myriad other processes (a trivial example being double diffractive processes leading to a pair of central jets with large rapidity gaps) and, hence, would have been long discovered.
In other words, if such an excess were to be observed, a single resonance hypothesis would be immediately ruled out. If it were to be then interpreted as an overlap of multiple resonances, two issues would need to be considered. The size of the excess imposes restrictions in the plane of the individual total widths and the diphoton branching fraction. On the other hand, a conservative estimate of the number of peaks can be obtained by comparing the width of the excess to the invariant mass resolution operative at such masses. And, finally, the very shape of the excess would dictate that the couplings of the said resonances must conform to a very specific pattern. To reproduce the features in a generic (i.e., non-clockwork) model (or even in multi-axion scenarios involving a new $N$-flavor confining sector Alexander et al. (2023, 2024)) would, thus, require many apparently ad hoc assumptions to be made. On similar grounds, it can also be argued that a broad resonance signal — with a finite bandwidth that is comparable with the characteristic mass scale of the resonance $\sim m$—is incompatible with, say, the Kaluza-Klein spectra of extra-dimensional scenarios consisting, in principle, of an infinite number of states (upto a natural cutoff of the theory), for then the distribution of resonances would extend over a large mass range and truncate only when the cross-sections start to fall near the kinematic limits of the experiment.
It is perhaps worth reiterating at this point that the analysis strategy employed here is complementary to the collider study of the clockwork graviton model carried out in refs.Giudice et al. (2018); Beauchesne and Kats (2020). Of the different strategies investigated therein, a particularly interesting approach involves searching for periodic signals in various kinematic distributions. However, the said technique, while well-suited for the tower-like spectra typically encountered in clockwork models, seems to be more efficient when the resonances are experimentally resolvable. The analysis above, therefore, presents a categorically new perspective to look at clockwork spectra at colliders, especially for the case of light ALPs with tiny mass-splittings.
IV.2 Intermediate mass ALPs
With an intermediate mass range set by the parameters $m=35$ GeV and $f=1000$ GeV, the analysis particulars in this case are largely the same as described for benchmark I. With the ALP masses in this case spanning the range 3510535-105 GeV, a sizable fraction of the spectrum still falls in the low mass region for which we require the final state photons to be boosted due to reasons delineated for benchmark I. We, therefore, choose to adopt the same strategy as in benchmark I for the collider simulation and consider a signal consisting of CW axions produced in association with upto two jets (and the axions further decaying to two photons). Photons and jets are reconstructed using the same criteria as in the case of Benchmark point I (see Table 1).
IV.2.1 Benchmark Point II(a)
The $p_{T}$ distributions for the leading and the sub-leading photons thus obtained for the signal events alongwith the corresponding distributions for the jet $p_{T}$’s are shown in Fig.9 (top), assuming a bin size of 1 GeV. The distributions are similar in nature to what was obtained for benchmark point I, albeit with a shift in the photon distributions towards higher $p_{T}$ values due to the ALPs being heavier in this case. The attendant diphoton invariant mass distribution is shown in Fig.9 (bottom). As for the SM backgrounds, for the mass range $m_{\gamma\gamma}\in[10,80]$ GeV, once again we adopt the ATLAS analysis Aad et al. (2023a). For the $[80,110]$ GeV window, in the absence of an appropriate analysis [^26] The ATLAS search for diphoton resonances in the mass range $m_{\gamma\gamma}=66-110$ GeV Aad et al. (2024b) does provide a background estimation in the region of interest. However, the event selection criteria employed therein is different than that in the low mass analysis that we refer to, which makes it difficult to match the background events corresponding to the low and intermediate mass ranges in a meaningful way., we choose to overestimate the background, by holding it at the level of $m_{\gamma\gamma}=80$ GeV. Thus, our significance estimate would be somewhat conservative.
Figure 9: BP-II(a): event distributions, after applying acceptance cuts alone, vide. Table 1. (top) $p_{T}$ distributions for the leading photon ($\gamma_{1}$), sub-leading photon ($\gamma_{2}$), leading jet ($j_{1}$) and the sub-leading jet ($j_{2}$) and (bottom) the two-photon invariant mass.
Figure 10: BP-II(a): Diphoton invariant mass distribution after applying selection cuts.
Figure 11: Bin-wise significance for benchmark II(a) at the LHC for $\mathcal{L}=138\,\rm fb^{-1}$.
To obtain the signal diphoton distribution in the fiducial region, we employ the selection cuts mentioned in Table 2. As for benchmark I, the kinematic cuts substantially reduce the number of events in the signal distribution, shown in Fig.10. Given the characteristic mass-splittings of this benchmark, we see that a large portion of the spectrum (consisting of 28 ALPs) in the diphoton distribution could potentially be resolved at the detector. Now, a bin-wise significance estimation, under the conservative assumption of the background events beyond $m_{\gamma\gamma}\approx 80$ GeV, leads to a distribution as shown in Fig.11 with a cumulative significance $S_{c,138}=2.12\sigma$ over the signal region. While spectral techniques Giudice et al. (2018); Beauchesne and Kats (2020) may enhance sensitivity in scenarios where multiple resonances are moderately separated and experimentally resolvable, the present analysis adopts a more conservative, profile-based strategy. This approach focuses on the overall signal shape, which effectively manifests as a single broad structure due to small mass splittings and finite detector resolution.
IV.2.2 Benchmark Point II(b)
As the only modification we have in this case is the increase in the number of ALPs to $N=40$, we follow the same strategy for the analysis as before. The signal’s diphoton invariant mass distribution post acceptance cuts is shown in Fig.12. In contrast with the outcome for benchmark II(a), the sizes of the individual peaks have decreased, with the cross-sections scaling as $\sim|C_{nN}|^{2}/N$, and their smearing is now relatively more pronounced as a result of the smaller mass-splittings, both effected by the increase in the number of ALPs. The corresponding distribution in the fiducial region is presented in Fig.13, exhibiting, expectedly, the same signal shape as in Fig.12, although with a reduced number of events. The bin-wise significance estimate is shown in Fig.14 with the corresponding cumulative significance being $S_{c,138}=1.97\sigma$.
Clearly, the two cases lead to slightly different projections for the $300\,\rm fb^{-1}$ luminosity reach, viz. $S^{(a)}_{c,300}\approx 3.12\sigma$ and $S^{(b)}_{c,300}\approx 2.9\sigma$. Thus, it may not be an overestimate to say that, upon a proper treatment, if the background distribution is found to saturate or slowly fall beyond $m_{\gamma\gamma}\sim 80$ GeV, then benchmark II could also be potentially probed by the end of LHC’s Run 3 phase.
For either the intermediate or the heavy-mass cases, the individual peaks are, thus, expected to be visible. The presence of such a bunch of closely packed resonances, with the specific envelope, would immediately point towards a clockwork-like scenario [^27] Naively, an alternative could be a composite sector with multiple excitations. However, not only would such a compositeness scale be very surprising, the abrupt ending of the excitations as well as the shape of the envelope would be hard to explain..
Figure 12: BP-II(b): Diphoton invariant mass distribution after applying acceptance cuts alone.
Figure 13: BP-II(b): Diphoton invariant mass distribution after applying selection cuts.
Figure 14: Bin-wise significance for benchmark II(b) at the LHC for $\mathcal{L}=138\,\rm fb^{-1}$.
IV.3 Heavy ALPs
For ALPs that are significantly heavier than considered hitherto, the decay photons would be expected to carry sufficient energy (and $p_{T}$) for the event to be triggered even without any additional jet. An example is afforded by the aforementioned Benchmark III, wherein the ALP masses span the range $\sim 150-450$ GeV. However, although extra jets are not needed for triggering, events with such jets do contribiute significantly to the signal cross sections. Hence, we define the signal as a semi-inclusive one composed of a pair of energetic photons with upto two additional jets [^28] The inclusion of events with three or more jets does not improve the signal to noise ratio.. The corresponding $K$-factors, for this mass range, are taken from refs.Ahmed et al. (2016); Boughezal et al. (2015).
Channel Acceptance Cuts
Photon identification
$pp\rightarrow a_{n}\,$ $\Delta R=0.4$, $p_{T}>0.4$ GeV, $p_{T}^{\text{ratio}}(\gamma)<0.12$
$a_{n}\rightarrow\gamma\gamma$ Jet identification
$\Delta R=0.4$ (anti-$k_{T}$), $p_{T}^{j}>20$ GeV, $|\eta_{j}|<2.5$
Isolation
$\Delta R(\gamma,\gamma)>0.4$, $\Delta R(\gamma,j)>0.4$, $\Delta R(j,j)>0.7$
$N_{\gamma}\geq 2$
Table 3: Acceptance cuts for the final state objects in benchmark III Aad et al. (2021).
The background, thus, would receive contributions from essentially the same channels as assumed in the low mass case. The signal profile being different, we, though, would need to adequately tune the assessment, and, in this, we are guided by the ATLAS diphoton analysis Aad et al. (2021) germane to this mass range. Post showering and hadronization, using Pythia8, we take for the identification and isolation of photons in the Delphes detector simulation a cone size of $\Delta R=0.4$ around the photon candidate. The requirement on the parameter $p_{T}^{\rm ratio}$, as defined earlier, is now reset to $p_{T}^{\rm ratio}<0.12$. The complete photon and jet identification criteria, as adapted from the ATLAS analysis ref.Aad et al. (2021), are displayed in Table 3. The $p_{T}$ distributions of the photons and the jets for the generated signal [^29] In contrast with benchmarks I and II, events without jets in the final state also contribute significantly in this case and, therefore, $N(2j)<N(1j)<N_{Tot}$. events are shown in Fig.15 (top), adopting a bin size of $16$ GeV as in the ATLAS searchAad et al. (2021).
Figure 15: BP-III: event distributions, after applying acceptance cuts alone, vide. Table 3. (top) $p_{T}$ distributions for the leading photon ($\gamma_{1}$), sub-leading photon ($\gamma_{2}$), leading jet ($j_{1}$) and the sub-leading jet ($j_{2}$) and (bottom) the two-photon invariant mass.
Fig.15 (bottom) shows the corresponding diphoton invariant mass distribution. Taking cue from ref.Aad et al. (2021), we then apply the selection cuts mentioned in Table 4 in order to define the fiducial volume in the phase space.
Channel Event Selection Criteria
$N_{\gamma}=2$, $N_{j}\leq 2$
$pp\rightarrow a_{n}\,$ $|\eta_{\gamma}|<2.37$ (excluding barrel-to-endcap region $1.37<|\eta_{\gamma}|<1.52$),
$a_{n}\rightarrow\gamma\gamma$ $E_{T}(\gamma_{1})>0.3\,m_{\gamma\gamma},\,E_{T}(\gamma_{2})>0.25\,m_{\gamma\gamma}$
$p_{T}^{j}>20$ GeV, $|\eta_{j}|<2.5$
Table 4: Event selection cuts applied to form the fiducial signal regions for benchmark III Aad et al. (2021).
The resulting diphoton invariant mass distribution is shown in Fig.16, whereas Fig.17 shows the corresponding bin-wise significance (with the background profile adopted from ref.Aad et al. (2021)) over the diphoton signal region with a cumulative significance $S_{c,138}\approx 1.58\sigma$. For this case, the projected enhancement for a $300\,\rm fb^{-1}$ luminosity reach is $S_{c,300}\approx 2.33\sigma$. Thus, in contrast to benchmarks I and II, the high mass scenario will not be readily accessible in the LHC’s Run 3. It is perhaps worth speculating, though, that benchmark III, with heavy ALPs, could potentially reach the discovery threshold during the forthcoming high luminosity phase of the LHC with a projected luminosity enhancement by a factor of nearly 20 times the current value.
Figure 16: BP-III—Diphoton invariant mass distribution after applying selection cuts.
Figure 17: Bin-wise significance for benchmark III at the LHC for $\mathcal{L}=138\,\rm fb^{-1}$.
V The Vector-like quarks and the heavy scalars
Beyond the low energy spectrum of the pseudoscalars, the model, as described in section II, also contains a heavy quark with the $SU(3)_{c}\times SU(2)_{L}\times U(1)_{Y}$ charge assignment $(3,1,2/3)$ as well as $N+1$ heavy radial scalar singlets (the partners of the pseudoscalars). Although not germane to the main objectives of our analysis, it is worth outlining here the dynamics of both the VLQ and the heavy scalars so as to establish the consistency of the model, especially in view of the fact that the assumed mass scales (characterised by the SSB scale $f$) of these heavy particles are, in principle, accessible at the LHC.
Considering that $m^{2}\ll f^{2}$ in all the benchmarks, the off-diagonal quadratic terms for the radial scalars in the Lagrangian (eq.(4)) are hierarchically smaller than the diagonal mass terms governed by the coupling $\lambda$, which is presumed to assume $\mathcal{O}(1)$ values. Consequently, all the radial scalars are nearly degenerate with masses $\sim\sqrt{2\lambda}f$. To simplify matters, we assume the other new quartic coupling, viz. $\lambda_{\Phi H}$, is also small, thereby automatically relaxing constraints from the stability of the SM Higgs potential, triviality, etc. Similar to the pseudoscalars, these heavy scalars have effective interactions with the gluons, photons and the EW vector bosons with the respective couplings given by $g_{\pi VV}/\xi$, where $g_{\pi VV}$ represent the pseudoscalar couplings as shown in eq.(17).
As stated in the discussion following eq.(8), we assume, for the sake of simplicity, that the singlet VLQ mixes predominantly with the SM top quark, thereby automatically suppressing FCNCs involving the first two generations of quarks. Post-EWSB, the two quarks mix and the corresponding transformations can be expressed (in the limit of neglecting the very small mixings with the first two generations) as
$\begin{split}\left(\Psi_{L}\quad u^{(3)}_{L}\right)^{T}&=U\,\left(T_{L}\quad t_{L}\right)^{T}\ ,\qquad{\rm and}\\ \left(\Psi_{R}\quad u^{(3)}_{R}\right)^{T}&=V\,\left(T_{R}\quad t_{R}\right)^{T}\ ,\end{split}$ (23)
where $T,t$ represent the mass eigenstates and $U,V$ are the (special unitary) mixing matrices. Since $\Psi_{R}$ and $u^{(3)}_{R}$ have identical quantum numbers, there is no gauge-mediated FCNC involving $T_{R}$ and $t_{R}$ and the weak gauge couplings of the VLQ-like state $T$ can be expressed as
$\begin{split}{\cal L}_{T-V}\displaystyle=\left(\frac{g}{\sqrt{2}}U_{tT}\right)W_{\mu}\,\bar{b}_{L}\gamma^{\mu}T_{L}\;+\left(\frac{g}{2c_{w}}U^{*}_{tt}U_{tT}\right)Z_{\mu}\,\bar{t}_{L}\gamma^{\mu}T_{L}+\displaystyle\frac{g}{c_{w}}\,Z_{\mu}\,\bar{T}\gamma^{\mu}\left[-\,\frac{2}{3}s_{w}^{2}+\frac{g}{2c_{w}}U^{*}_{Tt}U_{tT}P_{L}\right]T+\mbox{h.c.}\end{split}$ (24)
where $P_{L,R}$ are the usual chiral projection operators. On the other hand, the FCNCs involving the scalars, as derived from eq.(10), are given by
$\begin{split}\mathcal{L}_{T-\phi}&=-\frac{1}{\sqrt{2}}\bar{t}_{L}\left[\left(\lambda_{\Psi}U^{*}_{tT}V_{TT}+y^{\prime}_{\Psi}U^{*}_{tT}V_{tT}\right)\phi_{N}+\left(\lambda_{h}U^{*}_{tt}V_{tT}+y_{\Psi}U^{*}_{tt}V_{TT}\right)h\right]T_{R}\\ &-\frac{1}{\sqrt{2}}\bar{T}_{L}\left[\left(\lambda_{\Psi}U^{*}_{TT}V_{Tt}+y^{\prime}_{\Psi}U^{*}_{TT}V_{tt}\right)\phi_{N}+\left(\lambda_{h}U^{*}_{Tt}V_{tt}+y_{\Psi}U^{*}_{Tt}V_{Tt}\right)h\right]t_{R}+\mbox{h.c.}\ ,\end{split}$ (25)
where $\lambda_{h}$ denotes the top-Higgs Yukawa coupling. To obtain the pseudoscalar FCNCs, it is perhaps the most convenient to make use of the shift symmetry and rescale the field $\Psi_{L}\to e^{i\xi\pi_{N}/f}\Psi_{L}$ in eq.(10). This, in turn, results in a pseudovector FCNC term in the full Lagrangian through the kinetic term of the field $\Psi_{L}$, namely,
$\mathcal{L}_{T-\pi}=\frac{\partial_{\mu}\pi_{N}}{f}\,\bar{T}_{L}\gamma^{\mu}\left(U^{*}_{TT}U_{Tt}\right)t_{L}+\mbox{h.c.},$ (26)
with $\pi_{N}$ being expressed in terms of the mass eigenstates $a_{n}$ through $\pi_{N}=\sum_{k}C^{-1}_{Nk}a_{k}$ with the matrix $C$ as defined in eq.(6). For $f\gtrsim 1000$ GeV and $m_{T}\sim\lambda_{\Psi}f/\sqrt{2}\,>(m_{W,Z,h,\phi}+m_{t})$, the primary decay channels for the VLQ are clearly $T\to b\,W$, $T\to t\,Z$, $T\to t\,h$, $T\to t\,\phi_{N}$ (if kinematically allowed) and $T\to t\,a_{n}$ (see appendix B for the decay width expressions Atre et al. (2011)). Note that the decays to the new scalars and pseudoscalars appear in addition to the conventional VLQ decay modes to the SM particles and, therefore, it is worth comparing the branching fractions pertaining to the different channels. As this needs specifying a few further parameters, we begin by doing so for each of the benchmark points:
BP-I and II: We assume $\lambda_{\Psi}=2.2$, $y_{\Psi}=\epsilon y^{\prime}_{\Psi}$, $\lambda=1.8$ and $y^{\prime}_{\Psi}\lesssim 0.1$ with $\epsilon=0.1$. This choice of $\lambda_{\Psi}$ is motivated by the experimental lower bound on the VLQ mass ($m_{T}\gtrsim 1500$ GeV) to be discussed below. However, as can be expected from such a large value of the coupling, the evolution is fast and an examination of the three-loop renormalization group equations Buttazzo et al. (2013) leads to the conclusion that the theory becomes a very strongly coupled one at a scale $\mu\sim 10$ TeV, signalling that new physics must take over well before this scale [^30] A cutoff near $10$ TeV also serves to mitigate probable issues related to naturalness in the extended scalar sector, exhibiting only a little hierarchy problem.. Given that the clockwork model is not a UV-complete one, this, per se, may still be overlooked. However, we return to this point in the next section.
Note that the aforementioned collider limits assume that the VLQ decays within the detector. If, on the other hand, its decay length was large enough that it does not decay within the LHC detector [^31] This could happen if both $y^{\prime}_{\Psi}$ and $y_{\Psi}$ are tiny. The VLQ, nonetheless, has to decay on cosmological time scales so as not to violate the strong constraints on coloured and electrically charged dark matter candidates. Possibility of a bound state, made of stable neutral VLQs, being a DM candidate is discussed in De Luca et al. (2018)., it would hadronize and leave very distinctive signals. While such “R-hadrons” have been considered in the literature Baer et al. (1999); Arkani-Hamed and Dimopoulos (2005); Giudice and Romanino (2004); Kilian et al. (2005); Hewett et al. (2004); Arvanitaki et al. (2007); Fairbairn et al. (2007); Diaz-Cruz et al. (2007); Choudhury et al. (2008); Di Luzio et al. (2015), the corresponding experimental studies are not as exhaustive ATL (2019). The limits (typically ranging from $\sim 900$ GeV to $\sim 1250$ GeV) depend on the details of the hadronization model assumed. We refrain from considering this possibility in our analysis.
BP-III: The considerably larger value of the SSB scale $f$ allows us the luxury of choosing a relatively smaller value for the Yukawa coupling $\lambda_{\Psi}$, and we consider, instead, $\lambda_{\Psi}=1.5$, $\lambda=0.7$ and $y_{\Psi}=\epsilon y^{\prime}_{\Psi}$, $y^{\prime}_{\Psi}\lesssim 0.1$ with $\epsilon=0.1$. For such a choice, the running of $\lambda_{\Psi}$ is significantly slower and the strong coupling phase of the theory lies near $10^{8}$ GeV.
[…]
Branching Ratios
Channel SM BP-I & II BP-III
$T\to b\,W$ 0.5 0.44 0.47
$T\to t\,Z$ 0.25 0.21 0.23
$T\to t\,h$ 0.25 0.23 0.25
$T\to t\,a_{(\rm all)}$ 0.12 0.05
$m_{T}$ lower limit 1540 GeV Tumasyan et al. (2023); Alves et al. (2024) 1500 GeV $\approx$ 1540 GeV
Table 5: VLQ branching fractions for BP-I,$I$ and III (The SM column refers to the assumed branching fractions for VLQ searches Aad et al. (2023c); Tumasyan et al. (2023)). The bottom row displays the resulting lower bounds on $m_{T}$, as derived from the limits on the VLQ pair-production cross-section presented in ref.Tumasyan et al. (2023)
The constraints on the VLQ sector would, thus, come from direct observations at colliders [^32] Collider phenomenology of such KSVZ-VLQs have also been studied in refs.Ghosh et al. (2022); Ghosh and Konar (2023, 2024).. At the large hadron collider, the overwhelmingly leading production mechanism is the QCD-driven one [^33] With $U_{tT}$ not being large, single production is suppressed and the consequent bounds Aad et al. (2023c) are relaxed.. Once pair-produced, the $T$s would decay promptly. Table 5 lists the leading VLQ branching ratios for the three benchmark points [^34] For the chosen set of couplings, a decay of $T$ to radial scalars is kinematically forbidden. For any phenomenologically viable set of couplings, the branching fractions remain much smaller than those listed in the table.. It can be readily ascertained that the branching fractions have a very small dependence on the free parameter $y^{\prime}_{\Psi}$.
As the table shows, the branching fractions relevant to the standard search algorithms are not overly affected. Consequently, the derived limits are only slightly relaxed at best. On the other hand, it might be interesting to consider exotic channels such as $T\to t+a_{n}\to t+\gamma\gamma$. Hitherto (largely) unexplored, these might be of interest at future runs of the LHC.
VI Summary and Conclusion
We have examined the minimal QCD axion model within the clockwork paradigm and investigated the prospects of observing the massive ALPs that the model engenders at current and future hadron colliders. To this end, we employ a KSVZ-like setup with a heavy coloured $SU(2)_{L}$ singlet quark (as a top-partner) which couples to one end of a CW chain of $(N+1)$ complex scalars. Being chirally charged under the site-specific U(1)U(1)—which acts as an analogue of the PQ symmetry—this quark generates the requisite chiral anomaly for addressing the strong CP problem in a natural manner. The lightest pNGB in the CW sector identifies as the QCD axion while the accompanying heavier pNGBs in the spectrum behave as ALPs with a characteristic mass scale $\sim mq$. Through the chiral anomaly, these axions then couple to the gluons, the photon and the $Z$ boson, albeit with hierarchically [^35] The hierarchy is generated dynamically. different decay constants for the QCD axion and the ALPs, namely $f_{0}\sim q^{N}f$ and $f_{n>0}\sim f$, thanks to the CW mechanism. Being naturally consistent with the current experimental and observational constraints, the large suppression in the effective couplings of the light axion renders it practically invisible at high energy colliders (as is the case with all traditional QCD axions).
For a reasonable parameter configuration such as $q>1$, $N\sim\mathcal{O}(10)$, $m\gtrsim 10$ GeV and $f\sim 1$ TeV, what is more attractive instead is the production and detection of the ALPs, especially at the currently operating LHC and its future derivatives. To perform a quantitative analysis, we classify three benchmark scenarios according to the ALP masses in the range 103010-30 GeV (BP-I), 3510535-105 GeV (BP-II) and 150450150-450 GeV (BP-III). Such varied mass scales for the ALPs are enabled by the fact that the CW mechanism is based on the premise of localization in the theory space and, thus, is practically independent of the mass scales assumed in the theory. We performed an analysis of the expected signal profiles pertaining to the three benchmark points for the production of ALPs (with upto two additional jets) via gluon fusion and their subsequent decay to two photons at the LHC for an integrated luminosity of $138{\rm\,fb}^{-1}$. As anchor points for the analysis and for estimates of the pertinent background distributions, we referred to the ATLAS searches for diphoton resonances in the relevant invariant mass regions Aad et al. (2023a, 2021).
What stands out as particularly interesting is the case of the light ALPs where the number of particles $N$ that is consistent with the allowed window for the QCD axion’s decay constant also corresponds to very small mass-splittings among the ALPs. The resulting diphoton invariant mass distribution for the signal events turns out to be such that the individual resonances are smeared by the detector’s resolution (implemented in the form of a fast simulation of the ATLAS detector) and, thereby, overlap with each other with the full CW spectrum of light ALPs manifesting itself, instead, as a single but very broad resonance. It becomes immediately apparent, though, that a single such resonance is entirely untenable, and that such an excess must have resulted from the merging of several individual narrow resonances. On the other hand, treating the spectrum as a whole also offers the possibility of a better, more inclusive, estimation of the signal’s statistical significance over the background. For instance, the simplistic cumulative estimation defined in this work gives a sizable significance value of $S_{c}\sim 3.4\sigma$ (for an integrated luminosity of only $138~{\rm fb}^{-1}$) in BP-I, even though the significance estimates of the individual bins in the signal distribution are comparatively small. Moreover, when scaled by the projected luminosity reach of the ongoing Run 3 phase of the LHC, viz. $\sim 300{\rm\,fb}^{-1}$, the cumulative significance gets enhanced to $S_{c}\sim 5\sigma$. Of course, a dedicated search for such signals would call for a more detailed statistical analysis that is tailored to the fact that the individual resonances of the CW spectrum are characteristically correlated.
As for the heavy ALPs, e.g. those encountered in BP-III, the cumulative significance hovers just below $3\sigma$ even with $300{\rm\,fb}^{-1}$ of data. It should be borne in mind, though, that our analysis and projections are strictly based on the published ATLAS data. It is conceivable that a more refined analysis would serve to improve the sensitivity. What is exciting, though, is the prospect of probing a wide window of the CW ALPs at the HL-LHC with a planned luminosity reach of $\sim 3000{\rm\,fb}^{-1}$, where larger portions of the diphoton invariant mass profile could start to become apparent — the structure of the axion iceberg revealing itself — and even scenarios like BP-III could possibly surpass the discovery threshold. We note that the kind of multi-ALP signals discussed here possess a very broad and distinct quasi-periodic profile which cannot be reproduced in models with a single heavy ALP or in generic multi-scalar scenarios.
A key issue in our analysis has been the opposing pulls on the SSB scale $f$ from the need to increase the effective ALP–gluon–gluon coupling (scaling as $f^{-1}$) on the one hand and the lower bound on the VLQ mass (scaling as $f$) on the other. While modifying the number of sites $N$, or the exponent parameter $q$ does make a difference, the effect is a muted one (especially given that renormalizability of the low energy theory, as well as vacuum stability, requires that $q\leq 3$). However, given that the light ALP solutions presented here (benchmark points I & II) indicate that further new physics must exist by $\sim 10$ TeV scale, such a restriction might seem unwarranted and $q>3$ may be considered. However, since the ALP couplings to the gluons and photon have a dependence on $q$ which appears only through the elements of the CW transformation matrix, namely $C_{(n>0)N}$, the corresponding gain is marginal. It can be similarly argued that a change in the position of the VLQ coupling over the CW lattice (i.e., to assume $j<N$) also does not lead to a substantial gain.
Naively, a second possibility would be to postulate additional VLQs such that the effective $a_{n}gg$ coupling may be raised by way of all the VLQs contributing in the loop diagram. However, the corresponding change in the axion potential would putatively lead to the formation of stable domain walls/cosmic strings in the early Universe and is phenomenologically disfavoured. While one might attempt to mitigate the issue by invoking some of the solutions proposed in the literature, e.g. by assuming a soft breaking of the discrete symmetry or an inflation-induced statistical bias in the axion potential, the model would, nevertheless, be severely constrained due to the inherent drawbacks of such solutions Beyer and Sarkar (2023). Consider, instead, the introduction of one (or more) vector-like leptons (VLLs). This does not contribute to the $a_{n}gg$ couplings (thereby evading the domain wall constraints), but does contribute to the $a_{n}\gamma\gamma$ couplings, which now get enhanced by a factor of
$1+\frac{1}{3Y_{\Psi}^{2}}\sum N_{L}Q_{L}^{2}$
where $N_{L}$ is the number of VLLs of charge $Q_{L}$. Even the simplest choice of a single VLL of charge Q=1Q=-1 enhances the coupling by a factor of $1.75$, thereby allowing for a similar enhancement in $f$ without changing the rates. This, in turn, allows for a smaller Yukawa $\lambda_{\Psi}$, and postponement of the strong coupling phase until a cutoff scale $\mu\gtrsim 10^{7}$ GeV. As for the VLL phenomenology, the LHC limits are understandably much weaker. For it to decay, it must have at least a small Yukawa coupling with a SM lepton, and postulating this to be the $\tau$ would not only escape low-energy constraints, but also allow for intriguing signals at the LHC.
An entirely different mechanism of raising $f$ without suppressing the signal is afforded by the choice of the $U(1)_{N}$ charge $\xi$ for the $\Psi_{L}$ and $\Phi_{N}$. Increasing $\xi$ and $f$ proportionately keeps $f_{\rm eff}$ (and, therefore, the $a_{n}gg$ coupling) unchanged.
Whatever such mechanisms may be, it is quite apparent that the clockwork-axion-ALP scenario would be manifested at the LHC in terms of interesting signals, perhaps the most interesting being the case of the possibility of a strong coupling regime.
In addition to diphoton channels, several other prospective final states could enhance our search efforts at the collider, offering complementary avenues to test the model. Of particular interest are processes involving the heavy VLQ or the radial scalars which, while lying below the current LHC sensitivities, can be probed at the upcoming and proposed high-intensity and high-energy upgrades. For instance, as previously mentioned, the exotic decay of the VLQ to ALPs in association with a SM quark, e.g. $T\to t+a_{n}(\to\gamma\gamma)$, presents an interesting channel to study. Processes mediated by the heavy radial scalars $\phi_{n}$ such as $pp\to\phi_{n}\to a_{n}(\to 2\gamma/2j)a_{n}(\to 2\gamma/2j)$ are also worth exploring, more so because of the recent interest in the search for boosted diphotons Hayrapetyan et al. (2024). Another noteworthy process is the axion mediated production of a $Z$ boson in association with a photon, with the $Z$ subsequently decaying into dijets $(pp\to a_{n}\to Z(\to jj)\,\gamma)$. While the cross section for this process is low for the chosen benchmark points in the parameter space and the background is substantial, it remains a promising channel for probing the model at future colliders, particularly for benchmark point III.
This apart, there also exist other intriguing channels, the nonstandard decay $Z\to a_{n}\gamma$ being one interesting example. Although the process is hardly relevant for an investigation at the LHC, given the strong limits placed on the $Z$ exotic decay fractions at LEP, it could be well within the projected sensitivities of future lepton colliders, e.g. the proposed ILC running in the Giga-$Z$ mode Steinberg and Wells (2021). Furthermore, while the minimal setup considered in this work does not engender coupling of the ALPs to the light quarks or leptons even at the one-loop level, one could envision extending or modifying the model to include mixing of the VLQ with the light quarks and introduce additional VLLs that mix with the SM leptons in order to induce ALP couplings with the light SM fermions. In that case, several interesting channels open up, most notable of them being $h\to a_{n}a_{n}\to b\bar{b}\,\mu\bar{\mu}$ (with both the ATLAS and CMS experiments having observed intriguing signals Aad et al. (2022); Sirunyan et al. (2019) in this channel) and $h\to Z\,a_{n},\,a_{n}\to\,\mbox{hadrons}$ Schmieden (2021). This would also have interesting ramifications for flavour physics, including tests for lepton flavour universality in $B$-decays.
A complementary test of the model can also be found in the context of the direct detection of the light axion. A quick inspection shows that the masses and the effective photon couplings of the QCD axion in BP I and II(a) are close to the current sensitivities of several cavity haloscope experiments such as ADMX, CAPP, HAYSTAC, QUAX, etc. Thus, a collider probe of the characteristic multi-ALP signatures, which would largely constrain the parameters $m$ and $f_{\rm eff}$, alongwith the complementary probe of the light axion at the current or next generation of direct detection experiments, which is sensitive to the parameters $f_{\rm eff}$, $q$ and $N$, present a robust experimental platform to test the model in the near future.
Acknowledgements.
The authors thank Satyaki Bhattacharya, Tuhin S. Roy and Shilpi Jain for helpful discussions. D.C. acknowledges IoE, University of Delhi grant IoE/2024-25/12/FRP and grant no. CRG/2023/008234 of the ANRF, India. S.M. acknowledges research Grant No. CRG/2018/004889 of the SERB, India. T.S. acknowledges the support from the Dr. D.S. Kothari Postdoctoral fellowship scheme no. F.4-2/2006 (BSR)/PH/20-21/0163.
Appendices
Appendix A The axial anomaly
In the KSVZ model described in this work, the left and the right-handed projections of the heavy fermion $\Psi$ are chirally charged under the global symmetry $U(1)_{N}$. As is usually the case in KSVZ models, the chiral (or axial) $U(1)_{N}$ (analogous to $U(1)_{PQ}$ in the original setup) symmetry is anomalous which, in turn, results in the couplings of the pseudoscalar $\pi_{N}$ to gluons and the hypercharge boson given in eq.(11). This is most succinctly realized in the path integral formulation of the KSVZ theory, with the partition function given by
$Z[0]_{A}=\int\mathcal{D}\bar{\Psi}\mathcal{D}\Psi\exp{\left[i\int d^{4}x\,\mathcal{L}_{\Psi}\right]}.$ (27)
Here, the subscript $A$ denotes a fixed background with respect to the gauge fields and $\mathcal{L}_{\Psi}[\Psi,\bar{\Psi},D_{\mu}\Psi,D_{\mu}\bar{\Psi}]$ marks the full Lagrangian for $\Psi$ including the kinetic terms as well. Post SSB (at the scale $f$), the Yukawa term for $\Psi$ given in eq.(10) (ignoring the mixing terms with the SM fermions for brevity) can be rewritten in the convenient form
$\frac{\lambda_{\Psi}}{\sqrt{2}}(\phi+f)e^{i\xi\pi_{N}\gamma_{5}/f}\bar{\Psi}\Psi.$ (28)
The anomalous nature of $U(1)_{N}$ is apparent when we see that while the tree level action is invariant under the chiral transformation
$\Psi^{\prime}=U\Psi=e^{-i\xi\beta\gamma_{5}/(2f)}\Psi\,,\quad\pi_{N}\to\pi_{N}+\beta,$ (29)
where $\beta$ is a constant transformation parameter, the measure of the path integral is not. The measure would transform as,
$\mathcal{D}\bar{\Psi}\mathcal{D}\Psi\to\mathcal{J}^{-2}\mathcal{D}\bar{\Psi}\mathcal{D}\Psi\,$, (30)
where $\mathcal{J}$ is the Jacobian corresponding to the chiral transformation. The Jacobian, after appropriate regularisation, is given by Schwartz (2013); Bilal (2008)
$\mathcal{J}=\exp{\left[i\int d^{4}x\,\frac{1}{32\pi^{2}}\frac{\xi\,\beta}{f}\,\mbox{Tr}\{F^{\mu\nu}\tilde{F}_{\mu\nu}\}\right]}.$ (31)
where $F^{\mu\nu}\tilde{F}_{\mu\nu}\equiv F^{A\mu\nu}\tilde{F}^{B}_{\mu\nu}T^{A}T^{B}$ which stands for the gauge fields for all the local symmetries under which $\Psi$ is charged with $T^{A}$ being the respective generators. For brevity, we define $T^{A}$ such that they include the corresponding gauge couplings. Tr denotes trace over the gauge representation of the field $\Psi$. Therefore, the anomaly term which appears in the full Lagrangian is given by
$\mathcal{A}(x)=-\frac{\xi\,\beta}{16\pi^{2}f}\,\mbox{Tr}\{F^{\mu\nu}\tilde{F}_{\mu\nu}\}.$ (32)
Similarly, if we use the freedom to redefine the field $\Psi$ so as to absorb the pseudoscalar field through the chiral transformation
$\Psi\to e^{-i\xi\pi_{N}\gamma_{5}/(2f)}\Psi,$ (33)
we obtain the corresponding anomaly term in the Lagrangian given by
$\mathcal{A}(x)=-\frac{\xi\,\pi_{N}}{16\pi^{2}f}\mbox{Tr}\{F^{\mu\nu}\tilde{F}_{\mu\nu}\}.$ (34)
A straightforward evaluation of the trace in the preceding expression leads to the terms in eq.(11).
Appendix B VLQ decay widths
In the following we list the decay width expressions for the heavy fermion $\Psi$ with the labels $q$ and $V$ denoting the SM quarks $(b,t)$ and the EW gauge bosons $(W,Z)$ respectively.
$\begin{split}\Gamma(T\to qV)=&|g^{V}_{L}|^{2}\;\frac{\sqrt{m_{q}^{4}-2m_{q}^{2}\left(m_{T}^{2}+m_{V}^{2}\right)+\left(m_{T}^{2}-m_{V}^{2}\right)^{2}}}{32\pi\,m_{T}^{3}\,m_{V}^{2}}\Big{[}\left(m_{T}^{2}-m_{q}^{2}\right)^{2}+m_{V}^{2}\left(m_{T}^{2}+m_{q}^{2}\right)-2m_{V}^{4}\Big{]}\\[6.45831pt] \approx&\frac{|g^{V}_{L}|^{2}}{32\pi}\frac{m_{T}^{3}}{m_{V}^{2}}\quad(\mbox{for}\,\,m_{T}\gg m_{q,V})\quad.\end{split}$ (35)
Here, the couplings $g^{V}_{L}$ are given by
$g^{W}_{L}=\frac{g}{\sqrt{2}}U_{tT}\ ,\qquad\qquad g^{Z}_{L}=\frac{1}{2}\frac{g}{c_{w}}U^{*}_{tt}U_{tT}\ ,$ (36)
whereas the corresponding (off-diagonal) right-handed couplings vanish identically thanks to an analogue of the GIM mechanism.
For decays to scalars we ignore the small mixing between the SM Higgs and the heavy scalars $\phi_{n}$. The corresponding expressions are
$\begin{split}\Gamma(T\to t\,h)&=\frac{\sqrt{m_{h}^{4}-2m_{h}^{2}\left(m_{T}^{2}+m_{t}^{2}\right)+\left(m_{T}^{2}-m_{t}^{2}\right)^{2}}}{64\pi\,m_{T}^{3}}\\ &\times\left\{\left(|y^{h}_{Tt}|^{2}+|y^{h}_{tT}|^{2}\right)\left[m_{T}^{2}+m_{t}^{2}-m_{h}^{2}\right]+4\mbox{Re}\left(y^{h*}_{Tt}y^{h}_{tT}\right)m_{T}m_{t}\right\}\\[6.45831pt] &\approx\frac{|y^{h}_{Tt}|^{2}+|y^{h}_{tT}|^{2}}{64\pi}m_{T}\quad(\mbox{for}\,\,m_{T}\gg m_{t,h})\ ,\end{split}$ (37a)
$\begin{split}\Gamma(T\to t\,\phi_{N})&=\frac{\sqrt{m_{\phi_{N}}^{4}-2m_{\phi_{N}}^{2}\left(m_{T}^{2}+m_{t}^{2}\right)+\left(m_{T}^{2}-m_{t}^{2}\right)^{2}}}{64\pi\,m_{T}^{3}}\\ &\times\left\{\left(|y^{\Phi}_{Tt}|^{2}+|y^{\Phi}_{tT}|^{2}\right)\left[m_{T}^{2}+m_{t}^{2}-m_{\phi_{N}}^{2}\right]+4\mbox{Re}\left(y^{\Phi*}_{Tt}y^{\Phi}_{tT}\right)m_{T}m_{t}\right\}\\[6.45831pt] &\approx\frac{|y^{\Phi}_{Tt}|^{2}+|y^{\Phi}_{tT}|^{2}}{64\pi}m_{T}\left(1-\frac{m_{\phi_{N}}^{2}}{m_{T}^{2}}\right)^{2}\quad(\mbox{for}\,\,m_{T,\phi_{N}}\gg m_{t})\end{split}$ (37b)
and
$\begin{array}[]{rcl}\Gamma(T\to t\,a_{n})&=&\displaystyle\frac{\xi^{2}C_{nN}^{2}\sqrt{m_{n}^{4}-2m_{n}^{2}\left(m_{T}^{2}+m_{t}^{2}\right)+\left(m_{T}^{2}-m_{t}^{2}\right)^{2}}}{32\pi f^{2}\,m_{T}^{3}}\\[6.45831pt] &&\displaystyle\left(|U_{TT}|^{2}|U_{Tt}|^{2}\right)\left[\left(m_{T}^{2}-m_{t}^{2}\right)^{2}-\left(m_{T}^{2}+m_{t}^{2}\right)m_{n}^{2}\right]\\[10.76385pt] &\approx&\displaystyle\frac{\lambda_{\Psi}^{2}\,\xi^{2}\,C_{nN}^{2}|U_{TT}|^{2}|U_{Tt}|^{2}}{32\pi}\,m_{T}\qquad(\mbox{for}\,\,m_{T}\gg m_{t,a_{n}})\quad,\end{array}$ (37c)
where the effective Yukawa couplings are given by
$\begin{split}y^{h}_{Tt}&\approx\displaystyle\lambda_{h}U^{*}_{Tt}V_{tt}\ ,\\ y^{h}_{tT}&=\displaystyle\lambda_{h}U^{*}_{tt}V_{tT}+y_{\Psi}U^{*}_{tt}V_{TT}\ ,\\ y^{\Phi}_{tT}&\approx\displaystyle\lambda_{\Psi}U^{*}_{tT}V_{TT}\ ,\\ y^{\Phi}_{Tt}&=\displaystyle\lambda_{\Psi}U^{*}_{TT}V_{Tt}+y^{\prime}_{\Psi}U^{*}_{TT}V_{tt}\ .\end{split}$ (38)